### v1

# Opinion Clustering under the Friedkin-Johnsen Model: Agreement in Disagreement

Abstract
The convergence of opinions in the Friedkin-Johnsen (FJ) framework is well studied, but the topological conditions leading to opinion clustering remain less explored. To bridge this gap, we examine the role of topology in the emergence of opinion clusters within the network. The key contribution of the paper lies in the introduction of the notion of topologically prominent agents, referred to as Locally Topologically Persuasive (LTP) agents. Interestingly, each LTP agent is associated with a unique set of (non-influential) agents in its vicinity. Using them, we present conditions to obtain opinion clusters in the FJ framework in any arbitrarily connected digraph. A key advantage of the proposed result is that the resulting opinion clusters are independent of the edge weights and the stubbornness of the agents. Finally, we demonstrate using simulation results that, by suitably placing LTP agents, one can design networks that achieve any desired opinion clustering.
{IEEEkeywords}
Opinion dynamics; Friedkin-Johnsen model; Clustering of opinions.
1 INTRODUCTION
In recent decades, social media networks have become a major platform for advertising, conducting socio-political campaigns, and even spreading misinformation. This evolution has generated a keen interest among researchers to study the impact of social interactions on the formation of public opinion. Several mathematical models have been proposed, such as the DeGroot’s model [1], the FJ model [2], the Hegselman-Krause model [3], and the Biased assimilation model [4], that explain complex aspects of human interactions such as individual biases and homophily. Among these, the FJ model is popular because it captures a diverse range of emergent behaviours. Its popularity also stems from its analytical tractability and performance over large datasets.
The FJ model was proposed in [2] to justify the disagreement among closely interacting individuals. The opinions evolving under the FJ model converge to disagreement due to the presence of biased individuals, who are commonly known as the stubborn agents. In [2], the authors show that if each non-stubborn agent in the network has a directed path from a stubborn agent, the opinions converge asymptotically at a steady state. Further, they demonstrate that the final opinions depend on the initial opinions of only the stubborn agents and lie in the convex hull of these opinions. The authors in [5] introduced the notion of oblivious agents, who are non-stubborn and do not have a path from any stubborn agent. The algebraic and graph-theoretic conditions that ensure convergence in networks having oblivious agents were presented in [5] and [6], respectively. Additionally, in [6], the authors show that the final opinions depend on the initial opinions of both the stubborn agents and some oblivious agents.
In opinion dynamics, emergent behaviours such as consensus and opinion clustering are of keen interest to explain and achieve a desired social outcome. While consensus is being widely explored even in the FJ framework [7], this work focuses on opinion clustering. A network achieves opinion clustering when groups of two or more agents converge to the same opinion at a steady state. It represents swarm behaviour [8] and has applications in task distribution [9] and formation control [10]. The works [11, 12, 13] present graph-theoretic conditions under which opinion clustering occurs in continuous-time linear invariant systems. While it is known that disagreement occurs in the FJ framework, the conditions for opinion clustering are not well-explored. In [14], the authors present the graph-theoretic conditions to achieve opinion clustering in the FJ framework in networks with oblivious agents. However, the results apply only to those networks where each cycle contains only one stubborn agent.
Contrary to [14], our work aims to explain the emergence of opinion clusters in the FJ framework in arbitrarily connected digraphs. In opinion dynamics, an opinion cluster at steady state represents a group of individuals with a shared view on a topic. Our focus is specifically on how the network topology directs a certain group of agents to form an opinion cluster. With this, we now highlight our major contributions as follows:
  • We introduce a new type of agent called the Locally Topologically Persuasive (LTP) agent. We show that each LTP agent is associated with a unique set of non-influential agents in its vicinity.
  • We show that an LTP agent and the agents associated with it always form an opinion cluster under the FJ model. This result generalises the topology-based conditions presented in [14] to arbitrarily connected digraphs.
  • Importantly, we show that a non-stubborn LTP agent can steer the opinions of those associated with it to form an opinion cluster. Thereby, highlighting that even non-stubborn agents can impact the final opinions.
The paper has been organised as follows: Sec.2 introduces the notations and the relevant preliminaries. Sec. 3 motivates the analysis of the topological dependence of opinion clusters. Sec. 4 presents a framework to relate final opinions of any set of agents in the graph. We introduce the LTP agents and present the conditions for opinion clustering in Sec. 5. Finally, we conclude in Sec. 6 with insights into future directions.
2 Notations and Preliminaries
2.1 Notations
[…]
2.2 Graph Preliminaries
[…]
A walk is an ordered sequence of nodes such that each pair of consecutive nodes forms an edge in the graph. A cycle is a walk whose initial and final nodes coincide. If no nodes in a walk are repeated, it is a path. A graph is aperiodic if the lengths of all its cycles are coprime. An undirected graph is connected if a path exists between every pair of nodes. A directed graph (digraph) is a strongly connected graph if a directed path exists between every pair of nodes in the graph. A digraph is weakly connected if it is not strongly connected, but its undirected version is connected. A maximal subgraph of $\mathcal{G}$ which is strongly connected forms a strongly connected component (SCC). An SCC is an independent strongly connected component (iSCC) if every node in the SCC has all its in-neighbours within the same SCC.
2.3 Kron Reduction
Let $M\in\mathbb{R}^{p\times p}$ and $\alpha,\beta\subseteq[p]$ be index sets. Then $M[\alpha,\beta]$ is a submatrix of $M$ with rows indexed by $\alpha$ and columns indexed by $\beta$. The submatrix $M[\alpha]:=M[\alpha,\alpha]$ and $\alpha^{c}:=[p]\setminus\alpha$. If $M[\alpha^{c}]$ is non-singular, then the Schur complement of $M[\alpha^{c}]$ in $M$ is defined as:
$M/\alpha^{c}=M[\alpha]-M[\alpha,\alpha^{c}](M[\alpha^{c}])^{-1}M[\alpha^{c},\alpha]$ (1)
[…]
Lemma 1 ([16])
Consider a loopy Laplacian matrix $Q\in\mathbb{R}^{p}$ with $\alpha\subset[p]$. Let $\mathcal{G}(Q)$ be the digraph associated with $Q$.
  1. 1.
    $Q/\alpha^{c}$ is well defined if for each node $i\in\alpha^{c}$ there is a node $j\in\alpha$ such that a path $j\to i$ exists in $\mathcal{G}(Q)$. (Note that this condition is adapted to our framework because we consider $w_{ij}>0$ if edge $(j,i)\in\mathcal{E}$.)
  2. 2.
    If $Q$ is a loopless Laplacian matrix, then $Q/\alpha^{c}$ is also a loopless Laplacian matrix.
3 FJ model and emergent behaviours
Consider a network $\mathcal{G}$ of $n$ agents with vector $\mathbf{x}(k)=[x_{1}(k),...,x_{n}(k)]$ denoting the opinions of the agents at the $k^{th}$ instance. Under the FJ model, the opinions of agents evolve as follows:
$\mathbf{x}(k+1)=(I-\beta)W\mathbf{x}(k)+\beta\mathbf{x}(0)$ (2)
[…]
Suppose $\mathcal{G}$ is weakly connected and there is a non-stubborn agent who does not have a directed path from any stubborn agent. In [5], such agents are called oblivious agents. Without loss of generality, we can renumber the nodes in $\mathcal{G}$ such that the nodes $\{1,2,...,n_{o}\}$ are the oblivious agents and the rest are the non-oblivious, where $n_{o}\in[n]$. Additionally, we consider the oblivious agents that form an iSCC to be grouped together. Now, we can re-write the FJ-model (2) as:
$\mathbf{x}_{1}(k+1)$ $=W_{11}\mathbf{x}_{1}(k)$
$\mathbf{x}_{2}(k+1)$ $=(I-\bar{\beta})\big{(}W_{21}\mathbf{x}_{1}(k)+W_{22}\mathbf{x}_{2}(k)\big{)}+\bar{\beta}\mathbf{x}_{2}(0)$
where $\mathbf{x}_{1}(k)\in\mathbb{R}^{n_{o}}$ and $\mathbf{x}_{2}(k)\in\mathbb{R}^{(n-n_{o})}$ denote the opinions of oblivious and non-oblivious agents. The matrix $W$ gets partitioned as $W=\begin{bmatrix}W_{11}&0\\ W_{21}&W_{22}\end{bmatrix}$ and $\bar{\beta}\in\mathbb{R}^{(n-n_{o})\times(n-n_{o})}$ is a diagonal matrix with entries equal to the stubbornness of the non-oblivious agents.
Lemma 2 ([5, 6])
Consider a weakly connected network $\mathcal{G}$ with opinions evolving under the FJ model (2).
  • If $\mathcal{G}$ does not contain any oblivious agents, the final opinions $\mathbf{x}^{*}$ always converge to,
    $\mathbf{x}^{*}=(I-(I-\beta)W)^{-1}\beta\mathbf{x}(0)$
  • If $\mathcal{G}$ has oblivious agents, then the opinions converge at steady state only if each iSCC composed of oblivious agents is aperiodic. The final opinions converge to,
    $\mathbf{x}_{1}^{*}$ $=W_{11}^{*}\mathbf{x}_{1}(0)$
    $\mathbf{x}_{2}^{*}$ $=(I-(I-\bar{\beta})W_{22})^{-1}\big{(}(I-\bar{\beta})W_{21}W_{11}^{*}\mathbf{x}_{1}(0)$
    $+\bar{\beta}\mathbf{x}_{2}(0)\big{)}$
    where $W_{11}^{*}=\lim_{k\to\infty}W_{11}^{k}$.
Since $\mathbf{x}^{*}$ depends only on the initial opinions of the stubborn and the oblivious agents in iSCCs, these agents are called the influential agents. $\mathcal{I}$ denotes the set of all influential agents in $\mathcal{G}$. From Lemma 2, it follows that the impact of these influential agents on $\mathbf{x}^{*}$ depends on two factors: the network topology (captured by $W$) and the stubbornness of agents.
In general, the final opinions converge to different values, causing disagreement among agents. However, certain subgroups of agents can still have consensus amongst themselves and form opinion clusters in the final opinion. Naturally, the question arises how the formation of these opinion clusters depends on network topology and stubborn behaviour. The following example illustrates the effect of network topology on the opinion clusters.
Example 1
Consider the network $\mathcal{G}$ shown in Fig. 1(a). $\mathcal{G}$ has two stubborn agents: $2$ and $6$ with $\beta_{2}=0.3$, and $\beta_{6}=0.6$, and no oblivious agents. For initial opinions chosen from a uniform distribution on $[0,10]$, the opinions under the FJ model (2) evolve as shown in Fig. 1(b). The following opinion clusters form in the final opinion: agents $3,4$ and $5$ converge to $7.95$, agents $1,6$ converge to $8.85$ and agent $2$ converges to $6.83$. Next, we modify $\mathcal{G}$ by adding edge $(1,5)$ with weight $0.5$ (and reducing the weight of $(4,5)$ to $0.5$ to ensure $W$ remains row-stochastic) as shown in Fig.1(c). For the same stubbornness and the initial opinions, Fig. 1(d) shows that node $5$ forms a new cluster while the rest belong to the same opinion clusters.
(a) Network $\mathcal{G}$
(b) Evolution of opinions $\mathcal{G}$.
(c) Modified network $\mathcal{\hat{G}}$
(d) Evolution of opinions in $\mathcal{\hat{G}}$
Figure 1: Simple modification to $\mathcal{G}$ affects the opinion clusters in final opinion. Throughout the paper, we represent the stubborn agents and the non-stubborn agents by red nodes and orange nodes, respectively. The modified edges in $\mathcal{\hat{G}}$ are highlighted in red.
Example 1 demonstrates that the underlying topology has a significant impact on formation of opinion clusters in the network. Motivated by this, our objective is to establish the topology-based conditions that steer a subgroup of agents in the network to form an opinion cluster in the FJ framework.
4 The Relation of the final opinions with Network topology
In this section, we present a framework to relate the final opinions of agents with the underlying network topology. Consider a digraph $\mathcal{G}$ consisting of $n$ agents with opinions evolving under the FJ model (2). Let the network have $m$ stubborn agents. The final opinions satisfy the equations: $\mathbf{x}^{*}=(I-\beta)W\mathbf{x}^{*}+\beta\mathbf{x}(0)$. We can rewrite the above-mentioned linear equations as follows:
$\underbrace{\begin{bmatrix}(I-(I-\beta)W)&-\eta&\\ 0&0&\end{bmatrix}}_{R}\ \underbrace{\begin{bmatrix}\mathbf{x}^{*}\\ \mathbf{x}_{s}(0)\end{bmatrix}}_{z}=\mathbb{0}$ (3)
where $\mathbf{x}_{s}=[x_{s_{1}}(0),...,x_{s_{m}}(0)]$ and $s_{1},s_{2},...,s_{m}\in[m]$ are the labels of the stubborn agents. The matrix $\eta=[\eta_{ij}]\in\mathbb{R}^{n\times m}$ has entry $\eta_{{i}j}=\beta_{i}$ if the stubborn agent $i$ is labelled as $s_{j}$, otherwise $\eta_{ij}=0$ for $i\in[n],j\in[m]$. Henceforth, we represent eqn. (3) as $Rz=\mathbb{0}$ where $R=[r_{i,j}]\in\mathbb{R}^{(n+m)\times(n+m)}$.
Remark 1
The following are the salient properties of $R$: (a) each row sum in $R$ is 0, (b) the diagonal entries are non-negative while the off-diagonal entries are non-positive. Thus, $R$ is a Laplacian matrix under the Defn. 6.3 in [17].
The digraph $\mathcal{G}(R)$ can be induced from $R$ if an edge $(i,j)$ exists in $\mathcal{G}(R)$ when $r_{ji}<0$. The associated graph $\mathcal{G}(R)$ has $n+m$ nodes. Being consistent with the indexing of nodes in $\mathcal{G}$, the first $n$ nodes in $\mathcal{G}(R)$ are associated with $x_{i}^{*}$ for $i\in[n]$ and the nodes $n+1$ to $n+m$ are associated with the initial states of the $m$ stubborn agents. Note that each node $n+i$ forms a source in $\mathcal{G}(R)$ for $i\in[m]$. The source $n+i$ has a single outgoing edge to the node associated with the final opinion of the corresponding stubborn agent. For example, in Fig. 2, the initial opinion of stubborn agent $2$ is associated with node $n+1$ in $\mathcal{G}(R)$, thus, $n+1$ only has the outgoing edge $(n+1,2)$.
Example 2
Consider the network $\mathcal{G}$ shown in Fig. 1(a) with stubborn agents $2$ and $6$. The network $\mathcal{G}(R)$ for the matrix $R$ derived from eqn. (3) is shown in Fig. 2. The nodes in $\mathcal{G}(R)$ that represent the initial opinion of the stubborn agents are highlighted in turquoise.
Figure 2: The network $\mathcal{G}(R)$ derived from $\mathcal{G}$.
Remark 2
Since $R$ is a Laplacian matrix, it can be reduced by Kron reduction for a suitable choice of the subset of nodes $\alpha$ (defined in Sec. 2.3). Consider a network $\mathcal{G}$ without any oblivious agents. For such a network, each node in $[n]$ has a path in $\mathcal{G}(R)$ from at least one source $\{n+1,...,n+m\}$. Thus, by Statement 1) in Lemma 1, the Schur complement $R/\alpha^{c}$ is well-defined for any $\alpha$ such that $\{n+1,...,n+m\}\subseteq\alpha\subseteq[n+m]$. If $\mathcal{G}$ has an oblivious agent, then the sources $\{n+1,...,n+m\}$ do not have a path to this agent. Thus, we define $\alpha$ to include the oblivious agents belonging to each iSCC along with sources $\{n+1,...,n+m\}$. For this choice of $\alpha$, each $i\in\alpha^{c}$ has a path in $\mathcal{G}(R)$ from a node $j\in\alpha$. Thus, Statement 1) in Lemma 1 holds.
Importantly, eliminating the states associated with nodes in $\alpha^{c}$ in eqn. (3), reduces it to
$R/\alpha^{c}.\begin{bmatrix}\mathbf{x}^{*}[\omega]&\mathbf{x}_{s}(0)]\end{bmatrix}^{T}=\mathbb{0}$ (4)
where $\omega=\alpha\setminus\{n+1,...,n+m\}$. Thus, the Kron reduction technique is useful in determining the relations between the final opinions of a subgroup of nodes in $\mathcal{G}$ and quantifying their dependence on the influential agents (stubborn agents and oblivious agents in iSCC). Using the Kron-reduction technique, we present topology-based conditions under which the agents in a subgroup have equal final opinions and form an opinion cluster.
5 Topology-based opinion clustering
Consider a weakly connected network $\mathcal{G}$ with opinions evolving under the FJ model (2). Let $q$ be a non-influential agent in $\mathcal{G}$. Then, it is simple to see that the final opinion $x_{q}^{*}$ depends on the initial opinions of only those influential agents that have a path to $q$. Let $\mathcal{I}_{q}\subseteq\mathcal{I}$ be the set of all influential agents that have a directed path to $q$. Now, we define an LTP agent in terms of paths from agents in $\mathcal{I}_{q}$ to $q$.
Definition 1
Consider an agent $p$ in network $\mathcal{G}$. If every path in $\mathcal{G}$ from each influential agent $s\in\mathcal{I}_{q}$ to $q$ (where $p\neq q$) traverses $p$, then $p$ is an LTP agent and $q$ is said to be persuaded by $p$. Note that an LTP agent itself is not persuaded by any other agent, i.e. all paths from agents in $\mathcal{I}_{p}$ to $p$ must not pass through any other agent.
The set of all non-influential agents persuaded by an LTP agent $p$ is denoted as $\mathcal{N}_{p}$. Since $p$ is an LTP agent and $q\in\mathcal{N}_{p}$, it implies that all possible walks from each agent in $\mathcal{I}_{q}$ to $q$ also traverse $p$. An LTP agent can be an influential agent (i.e. a stubborn agent or an oblivious agent in an iSCC) or a non-influential agent. An influential agent is an LTP agent only if each path from the remaining influential agents in $\mathcal{I}_{q}$ to $q$ traverses $p$. The following example illustrates the identification of an LTP agent and the agent(s) it persuades.
Example 3
Consider the network $\mathcal{G}$ shown in Fig. 1(a). Since $\mathcal{G}$ is strongly connected, only the stubborn agents $2$ and $6$ are influential and have a path to all agents. Fig 3(a) demonstrates that each path from $2$ to $4$ (and $5$) in $\mathcal{G}$ traverses through $3$. Moreover, the paths from $6$ to $4$ (and $5$) also traverse through $3$ as displayed in Fig. 3(b). Since all paths from $2$ and $6$ to $3$ do not traverse any common node. Thus, $3$ is an LTP agent and $\mathcal{N}_{3}=\{4,5\}$. Similarly, $6$ is also an LTP agent and $\mathcal{N}_{6}=\{1\}$.
(a) Paths from $2$ to $4$ (and $5$) traverse 33
(b) Paths from $6$ to $4$ (and $5$) traverse 33
Figure 3: Node 33 is an LTP agent and it persuades 44 and $5$ as each path from the stubborn agents traverse $3$.
Suppose a network $\mathcal{G}$ has multiple LTP agents. The following question naturally arises: Can an agent be persuaded by two different LTP agents? The following Lemma presents an answer to this question.
Lemma 3
Consider a network $\mathcal{G}$ and let $\mathcal{L}$ denote the set of LTP agents in $\mathcal{G}$. Then, $\mathcal{N}_{i}\cap\mathcal{N}_{j}=\emptyset$ for any distinct $i,j\in\mathcal{L}$.
Proof 5.1.
Suppose the contrary is true and there exists a node $q\in\mathcal{N}_{i}\cap\mathcal{N}_{j}$. Then, by definition, each path in $\mathcal{G}$ from every influential agent in $\mathcal{I}_{q}$ to $q$ must traverse both $i$ and $j$. The following scenarios can occur:
  1. C1
    Every path from $s$ to $q$ traverses $i$ before $j$ for all $s\in\mathcal{I}_{q}$.
  2. C2
    Every path from $s$ to $q$ traverses $j$ before $i$ for all $s\in\mathcal{I}_{q}$.
  3. C3
    Consider $s,u\in\mathcal{I}_{q}$. There exists a path from $s$ to $q$ that traverses $i$ before $j$, and there exists another path from $u$ to $q$ that traverses $j$ before $i$. Here, $s$ and $u$ can be the same node as well.
Under C1, every path from an agent in $\mathcal{I}_{q}$ to $j$ also traverses $i$. Since $j$ has a path to $q$, it follows that $\mathcal{I}_{j}\subseteq\mathcal{I}_{q}$, and hence every path from each node in $\mathcal{I}_{j}$ to $j$ traverses $i$. Thus $j\in\mathcal{N}_{i}$, so $j$ cannot be an LTP agent. Equivalently, under C2, we obtain $i\in\mathcal{N}_{j}$, so $i$ is not an LTP agent. Hence, both C1 and C2 result in contradiction.
Under C3, there exists a path $P_{1}$ from $s$ to $q$ that traverses $i$ before $j$ and another path $P_{2}$ from $u$ to $q$ that traverses $j$ before $i$. We can represent $P_{1}$ as $s\to i\to j\to q$ and $P_{2}$ as $u\to j\to i\to q$. By definition, a path does not contain repeated nodes. It follows from $P_{1}$ that a path $j\to q$ exists that does not traverse $i$. Further, we know from $P_{2}$ that a path from $u\to j$ does not traverse $i$. Therefore, a path from $u\to q$ exists that does not traverse $i$. Consequently, $q\notin\mathcal{N}_{i}$ which contradicts our assumption that $q\in\mathcal{N}_{i}\cap\mathcal{N}_{j}$. Similarly, we can show that a path from $s\to q$ exists that does not traverse $j$. Since all of the conditions: C1, C2 and C3, lead to a contradiction, it means that the assumption $q\in\mathcal{N}_{i}\cap\mathcal{N}_{j}$ does not hold.
Lemma 3 establishes that an LTP agent and the agents it persuades form a disjoint group in the network. However, there can be certain agents who are neither an LTP agent nor are persuaded by any LTP agent. Excluding them, we can uniquely partition the remaining agents into disjoint groups, with each group consisting of one LTP agent and its persuaded agents. The following results presents the relation between the final opinions of agents within each such group.
Lemma 5.2.
Consider a weakly connected network $\mathcal{G}$ with matrix $R$ derived using eqn. (3). Let $p$ be an LTP agent and $q\in\mathcal{N}_{p}$. If $\alpha$ is the union of set $\{p,q,n+1,....,n+m\}$ and all the oblivious agents in iSCC in $\mathcal{G}$ and $\alpha^{c}=[n+m]\setminus\alpha$ such that $\alpha^{c}\neq\phi$. Then,
$(R[\alpha^{c}])^{-1}=\sum_{k=0}^{\infty}\big{(}(I-\beta[\alpha^{c}])W[\alpha^{c}]\big{)}^{k}$ (5)
Proof 5.3.
By definition of $\alpha$, we get $\alpha^{c}\subset[n]$. From the construction of $R$, it follows that $R[\alpha^{c}]=I-(I-\beta[\alpha^{c}])W[\alpha^{c}]$. If $\rho((I-\beta[\alpha^{c}])W[\alpha^{c}])<1$, then identity (5) follows by Neumann series. Let $M=(I-\beta[\alpha^{c}])W[\alpha^{c}]$. To prove that $\rho(M)<1$, it suffices to show that (I) $M$ is row substochastic, (II) for each $i^{th}$ row in $M$ with row-sum is $1$ there exists a $j^{th}$ row with row-sum less than $1$ such that there is a path from $j$ to $i$ in the associated digraph $\mathcal{G}(M)$. Then, by Theorem 6.37 in [18], $\rho(M)<1$.
Step-I: $M$ can be row-substochastic because: (i) a node $i\in\alpha^{c}$ whose in-neighbour is $p,q$ or an oblivious agent in iSCC satisfies $\sum_{j\in\alpha^{c}}w_{ij}<1$, (ii) each stubborn agent $i$ has the corresponding row-sum at most equal to $(1-\beta_{i})<1$. In an arbitrary digraph, the following scenarios may occur:
  • $h\in\alpha^{c}$ has an oblivious agent: Due to weak connectivity of $\mathcal{G}$, an oblivious agent in iSCC exists in $\alpha$ that has a path to $h$, implying that (i) holds.
  • All agents in $\alpha^{c}$ are non-oblivious and a stubborn agent exists in $\alpha^{c}$: in this case, (ii) holds for the stubborn agent.
  • All agents in $\alpha^{c}$ are non-oblivious and non-stubborn: Since $q$ cannot be stubborn, node $p$ is the only stubborn agent in $\mathcal{G}$. Thus, at least one node exists in $\alpha^{c}$ that has either $p$ or $q$ as in-neighbour, implying that (i) holds.
Thus, there always exists a node in $\alpha^{c}$ that satisfies either (i) or (ii). It implies that $M$ is row substochastic.
Step-II: As discussed in Step-I, each oblivious agent has a path from an oblivious agent in iSCC in $\mathcal{G}$. It follows from (i) that this path traverses an oblivious agent in $\alpha^{c}$ whose row-sum in $M$ is less than $1$. A special case could occur if each path in $\mathcal{G}$ from the oblivious agents in iSCC to an oblivious agent $\tilde{j}\in\alpha^{c}$ either traverses $p$ or $q$. In this case, the in-neighbours of $p$ or $q$ in $\alpha^{c}$ who have row-sum in $M$ less than $1$ lie on this path. Hence, every oblivious agent in $\alpha^{c}$ with row sum $1$ has a path in $\mathcal{G}(M)$ from a node with row sum <1<1. A similar argument applies to the non-oblivious agents as well. Thus, statement II also holds and $\rho(M)<1$.
Theorem 1.
Consider a network $\mathcal{G}$ of $n$ agents with $m$ stubborn agents and no oblivious agents. The opinions of agents in $\mathcal{G}$ are governed by the FJ model (2). If $p$ is an LTP agent, then opinions of $p$ and nodes in set $\mathcal{N}_{p}$ converge to the same final opinion and form an opinion cluster.
Proof 5.4.
Let $q\in\mathcal{N}_{p}$. To prove that the final opinion $x_{p}^{*}=x_{q}^{*}$, we reduce the matrix $R$ in eqn. (3) using the Kron reduction and examine the relation between $x_{p}^{*}$ and $x_{q}^{*}$ in the reduced set of linear equations (4). Since $\mathcal{G}$ does not have any oblivious agents, each non-stubborn agent has a path from at least one stubborn agent. From the discussion in Remark 2, we know that the Kron Reduction $R/\alpha^{c}$ is well-defined for $\alpha=\{p,q,n+1,...,n+m\}\supset\{n+1,...,n+m\}$. Next, we evaluate the Schur complement $R/\alpha^{c}$ given by eqn. (1).
The following structural properties of the associated network $\mathcal{G}(R)$ will be use in determining $R/\alpha^{c}$,
  1. 1.
    Each walk from $s\in\mathcal{I}_{q}$ to $q$ in $\mathcal{G}$ traverses $p$ for all $s\in\mathcal{I}_{q}$. Thus, by construction of $\mathcal{G}(R)$, each walk that from $n+i$ to $q$ in $\mathcal{G}(R)$ will traverse $p$ for all $i\in[m]$.
  2. 2.
    The nodes n+1,,n+mn+1,...,n+m in $\mathcal{G}(R)$ are sources.
Property 1) implies that $r_{q,n+i}=0$ for all $i\in[m]$ because node $n+i$ cannot be an in-neighbour of $q$ in $\mathcal{G}(R)$. Further, property 2) implies that the rows of $R$ indexed $\{n+1,...,n+m\}$ have all entries as zero. Thus,
$R[\alpha]=\begin{bmatrix}r_{p,p}&r_{p,q}&r_{p,n+1}&...&r_{p,n+m}\\ r_{q,p}&r_{q,q}&0&...&0&\\ \mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&...&\mathbb{0}_{m\times 1}\end{bmatrix}$.
Next, we consider $Y=R[\alpha,\alpha^{c}](R[\alpha^{c}])^{-1}R[\alpha^{c},\alpha]$. From Lemma 5.2, it follows that
$Y=R[\alpha,\alpha^{c}]\sum_{k=0}^{\infty}\big{(}(I-\beta[\alpha^{c}])W[\alpha^{c}]\big{)}^{k}R[\alpha^{c},\alpha]$ (6)
Let $F:[m+2]\to\alpha$ be a mapping such that $F(1)=p,F(2)=q$ and $F(j+2)=n+j$ for $j\in[m]$. This mapping relates the rows (and columns) indices of $Y$ (and $R/\alpha^{c}$) with the agents in set $\alpha$. Then, from eqn. (6), the entry $y_{ij}$ of $Y$ is non-zero only if there exists a walk in $\mathcal{G}(R)$ from $F(j)$ to $F(i)$ such that each node on this walk (except $F(i)$ and $F(j)$) belongs to $\alpha^{c}$. By property 1), $y_{2j}=0$ for all $j\in\{3,...,m+2\}$ because each walk from n+j2n+j-2 to $q$ in $\mathcal{G}(R)$ always passes through $p\in\alpha$. Consequently, $R/\alpha^{c}=R[\alpha]-Y$ has the following form:
$R/\alpha^{c}=\begin{bmatrix}r_{1,1}^{1}&r_{1,2}^{1}&r_{1,3}^{1}&...&r_{1,m+2}^{1}\\ r_{2,1}^{1}&r_{2,2}^{1}&0&...&0\\ \mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&...&\mathbb{0}_{m\times 1}\end{bmatrix}$ where $r_{i,j}^{1}$ gives the entry of matrix $R/\alpha^{c}$ for $i,j\in[m+2]$. From Lemma 1, we know that $R/\alpha^{c}$ is also a Laplacian matrix and the row-sum of each of its rows equals 0. Since the second row of $R/\alpha^{c}$ has only two non-zero entries, thus, $r_{2,1}^{1}=-r_{2,2}^{1}$. Moreover, by eliminating the states $\alpha^{c}$ in the steady state eqns. (3), we get the following reduced system:
$\begin{bmatrix}r_{1,1}^{1}&r_{1,2}^{1}&r_{1,3}^{1}&...&r_{1,m+2}^{1}\\ -r_{2,2}^{1}&r_{2,2}^{1}&0&...&0\\ \mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&...&\mathbb{0}_{m\times 1}\end{bmatrix}\begin{bmatrix}x_{p}^{*}\\ x_{q}^{*}\\ \mathbf{x}_{s}(0)\end{bmatrix}=\mathbb{0}$ (7)
Simplifying eqn. (7) shows that $x_{p}^{*}=x_{q}^{*}$. Since $q$ is any node in $\mathcal{N}_{p}$, this holds for all nodes in $\mathcal{N}_{p}$.
Theorem 1 establishes that each LTP agent $p$ and the agents in $\mathcal{N}_{p}$ collectively form an opinion cluster in the final opinion. Importantly, the formation of an opinion cluster depends on the existence of an LTP agent in a network, independent of the edge weights and the stubbornness of the stubborn agents. The Example 1 illustrates the result in Theorem 1 as the LTP agent $3$ forms an opinion cluster with agents in $\mathcal{N}_{3}=\{4,5\}$ and LTP agent $6$ forms another opinion cluster with agent in $\mathcal{N}_{6}=\{1\}$. Additionally, it explains the formation of a new opinion cluster for the modified graph $\mathcal{\hat{G}}$. This occurs because the addition of edge $(1,5)$ results in a path from $6$ to $5$ that does not traverse $3$. Thus, $5\notin\mathcal{N}_{3}$ and forms a new opinion cluster in Fig. 1(d).
Remark 5.5.
An LTP agent can be a stubborn or a non-stubborn agent. However, it is crucial in ensuring that the nodes in its close vicinity in $\mathcal{N}_{p}$ form an opinion cluster. Thus, even a non-influential LTP agent impacts the final opinion formation, making it topologically influential.
The following result extends Theorem 1 to the networks containing the oblivious agents.
Theorem 2.
Consider a network $\mathcal{G}$ of $n$ agents with $m$ stubborn agents and at least one oblivious agent. The opinions of agents in $\mathcal{G}$ are governed by the FJ model (2). If each iSCC in $\mathcal{G}$ composed of the oblivious agents is aperiodic, then final opinions of an LTP agent $p$ and nodes in set $\mathcal{N}_{p}$ form an opinion cluster.
The proof is along the same lines as Theorem 1 and is omitted for brevity. By definition, if an LTP agent $p$ is oblivious, then the agents in $\mathcal{N}_{p}$ are also oblivious. Consequently, it follows from Theorem 2 that a set of non-influential oblivious agents can also form an opinion cluster. Since the opinion of an oblivious agent evolves according to DeGroot’s model and is not affected by stubborn behaviour, an LTP agent ensures the formation of opinion clusters even in DeGroot’s framework. Note that since nodes in $\mathcal{N}_{q}$ are always non-influential, the desired opinion clusters can be formed in DeGroot’s framework only in weakly connected networks.
Remark 5.6.
In [14], a network with $m$ stubborn agents and no oblivious agents has $m$ opinion clusters in the final opinion if the agents can be partitioned into subgraphs such that each subgraph contains only one stubborn agent and has a spanning tree with the stubborn agent as its root. Additionally, only the stubborn agent in a subgraph can receive information from the other subgraphs. An equivalent result exists for networks with oblivious agents as well. Note that it is assumed that each cycle in the network contains only one stubborn agent.
The topological conditions in [14] imply that all the paths to a node in a subgraph from the rest of the stubborn agents traverse the stubborn agent to the subgraph in which the node belongs. Thus, this stubborn agent is effectively the LTP agent for the agents within a subgraph that collectively form an opinion cluster. Hence, Theorems 1 and 2 extend the results presented in [14] to any weakly connected digraph.
Using the topological conditions in Theorems 1 and 2, a network can be designed to obtain a predefined number of opinion clusters with a desired subset of agents forming them. The following example highlights this property.
Example 4
Consider the network $\mathcal{G}$ with stubborn agents $8$ and $10$ in Fig. 4(a). The final opinion is desired to form opinion clusters with each cluster containing the agents within the dashed boxes in Fig. 4(a). To achieve desired clustering, we design $\mathcal{G}$ such that each dashed box (except the iSCC composed of oblivious agents that already form an opinion cluster) has an LTP agent and the remaining agents are persuaded by it. In Fig. 4(a), the nodes with green boundaries denote the LTP agents. For initial conditions from a uniform distribution over $[0,10]$ and stubbornness values in $[0,1)$, the opinions of agents in $\mathcal{G}$ under the FJ model form the desired opinion clusters as shown in Fig. 4(b)
(a) Network $\mathcal{G}$
(b) Formation of desired opinion clusters
Figure 4: Suitable design of $\mathcal{G}$ for desired opinion clusters.
6 Conclusions
The FJ model incorporates the individual biases as stubborn behaviour in the averaging-based opinion models. Due to varied biases, even the individuals in a closely connected group hold diverse opinions. Even then, a subgroup of individuals can still reach consensus, thereby, forming an opinion cluster. This work examines the role of the underlying network topology in the emergence of these clusters. To begin with, we define a topologically special agent called an LTP agent, and a set of non-influential agents that it persuades. Every path from an influential agent (a stubborn agent or an oblivious agent in an iSCC) to the persuaded agent always traverses the LTP agent. Using Kron-reduction, we establish in Theorems 1 and 2 that the final opinions of an LTP agent and the set of agents it persuades are equal. Thus, they form an opinion cluster in any arbitrarily-connected digraph under the FJ framework. Interestingly, such LTP agents lead to the formation of opinion clusters even under the DeGroot’s model (see Theorem 2).
Some additional key insights from our results are: (i) The presence of LTP agents always results in opinion clusters, independent of the edge-weights in the network and the stubbornness of the agents. (ii) Notably, an LTP agent may or may not be influential, but still shapes the final opinions of the agents it persuades. (iii) The notion of LTP agents generalises the topology-based conditions for opinion clustering in [14] to any arbitrary digraphs. (iv) By suitably placing the LTP agents and designing the network topology, we achieve any desired opinion clustering as demonstrated in Example 4.
In future, we plan to examine the potential of LTP agents in desirably shaping the final opinions in real-world social networks.

### v2

@@
 A key advantage of the proposed result is that the resulting opinion clusters are independent of the edge weights and the stubbornness of the agents.
-Finally, we demonstrate using simulation results that, by suitably placing LTP agents, one can design networks that achieve any desired opinion clustering.
+Finally, we demonstrate that by suitably placing LTP agents, one can design networks that achieve any desired opinion clustering.
 {IEEEkeywords}
@@
 1 INTRODUCTION
-In recent decades, social media networks have become a major platform for advertising, conducting socio-political campaigns, and even spreading misinformation.
-This evolution has generated a keen interest among researchers to study the impact of social interactions on the formation of public opinion.
-Several mathematical models have been proposed, such as the DeGroot’s model [1], the FJ model [2], the Hegselman-Krause model [3], and the Biased assimilation model [4], that explain complex aspects of human interactions such as individual biases and homophily.
-Among these, the FJ model is popular because it captures a diverse range of emergent behaviours.
-Its popularity also stems from its analytical tractability and performance over large datasets.
-The FJ model was proposed in [2] to justify the disagreement among closely interacting individuals.
-The opinions evolving under the FJ model converge to disagreement due to the presence of biased individuals, who are commonly known as the stubborn agents.
-In [2], the authors show that if each non-stubborn agent in the network has a directed path from a stubborn agent, the opinions converge asymptotically at a steady state.
-Further, they demonstrate that the final opinions depend on the initial opinions of only the stubborn agents and lie in the convex hull of these opinions.
-The authors in [5] introduced the notion of oblivious agents, who are non-stubborn and do not have a path from any stubborn agent.
-The algebraic and graph-theoretic conditions that ensure convergence in networks having oblivious agents were presented in [5] and [6], respectively.
-Additionally, in [6], the authors show that the final opinions depend on the initial opinions of both the stubborn agents and some oblivious agents.
-In opinion dynamics, emergent behaviours such as consensus and opinion clustering are of keen interest to explain and achieve a desired social outcome.
-While consensus is being widely explored even in the FJ framework [7], this work focuses on opinion clustering.
-A network achieves opinion clustering when groups of two or more agents converge to the same opinion at a steady state.
-It represents swarm behaviour [8] and has applications in task distribution [9] and formation control [10].
-The works [11, 12, 13] present graph-theoretic conditions under which opinion clustering occurs in continuous-time linear invariant systems.
-While it is known that disagreement occurs in the FJ framework, the conditions for opinion clustering are not well-explored.
+The recent evolution of social media networks has generated a keen interest among researchers to study the impact of social interactions on public opinion.
+Several mathematical models such as the DeGroot’s model [1], the FJ model [2], the Hegselmann-Krause model [3], and the Biased assimilation model [4], explain complex aspects of human interactions such as individual biases and homophily.
+Among these, the FJ model is most analytically tractable and explains the frequently observed phenomenon of disagreement among closely interacting individuals [5].
+Moreover, it is widely popular because it is empirically validated on real-world datasets [6], [7].
+The opinions evolving under the FJ model converge to disagreement due to the presence of biased individuals, who are called stubborn agents.
+In [2], the authors show that if each non-stubborn agent in the network is reachable from a stubborn agent, the final opinions depend on the initial opinions of only the stubborn agents.
+Reference [8] introduces the notion of oblivious agents, who are non-stubborn and do not have a path from any stubborn agent.
+The algebraic and graph-theoretic conditions that ensure convergence in networks with oblivious agents are presented in [8] and [9], respectively.
+In opinion dynamics, emergent behaviours such as consensus, polarisation and opinion clustering are of interest to explain and achieve a desired social outcome.
+Several works [6, 10, 11, 12] have examined the FJ framework to reach these behaviours.
+In [6], the authors design stubbornness values that lead to behaviours such as consensus and polarisation.
+On the other hand, reference [10] presents topology-based conditions for graphs with two stubborn agents that result in polarisation.
+Furthermore, the authors in [11] examine the FJ model with randomised interactions and reveal that the opinions of non-stubborn agents can exhibit oscillations in presence of stubborn agents.
+In contrast, the present work focuses on opinion clustering, which is less-explored in the FJ framework.
+Opinion clustering occurs when a group of two or more agents converge to the same final opinion.
+It represents swarm behaviour and has applications in task distribution [13] and formation control [14].
+In the sociological framework, opinion clustering represents the formation of groups with aligned views.
+In this work, we examine the network topology towards achieving opinion clustering.
+Recent works [10, 11] in the FJ framework highlight the impact of topology in shaping emergent behaviours.
+More recently, the study of network topology is receiving attention to analyse the effects of recommender algorithms on opinion formation.
+The seminal works [15] and [16] contributed significantly by quantifying the influence of a stubborn agent using hitting probabilities of random walks over the network.
+These studies enabled the design of topology-based interventions (recommendations) for desirably shaping opinions [16, 17].
+In the present work, we further explore the relation between the network topology and final opinions and address the following problem: Under what topology-based conditions, does a set of agents converge to the same final opinion (form an opinion cluster)?
+We leverage this relationship to obtain desired opinion clustering.
 In [12], the authors present the graph-theoretic conditions to achieve opinion clustering in the FJ framework in networks with oblivious agents.
-However, the results apply only to those networks where each cycle contains only one stubborn agent.
+However, the result applies only to networks where each cycle contains at most one stubborn agent.
 Contrary to [12], our work aims to explain the emergence of opinion clusters in the FJ framework in arbitrarily connected digraphs.
-In opinion dynamics, an opinion cluster at steady state represents a group of individuals with a shared view on a topic.
-Our focus is specifically on how the network topology directs a certain group of agents to form an opinion cluster.
-With this, we now highlight our major contributions as follows:
-  • We introduce a new type of agent called the Locally Topologically Persuasive (LTP) agent. We show that each LTP agent is associated with a unique set of non-influential agents in its vicinity.
+With this, we now highlight our major contributions:
+  • We define the notion of the Locally Topologically Persuasive (LTP) agent. Each LTP agent is associated with a unique set of non-influential agents in its vicinity.
   • We show that an LTP agent and the agents associated with it always form an opinion cluster under the FJ model.
 This result generalises the topology-based conditions presented in [12] to arbitrarily connected digraphs.
-  • Importantly, we show that a non-stubborn LTP agent can steer the opinions of those associated with it to form an opinion cluster.
-Thereby, highlighting that even non-stubborn agents can impact the final opinions.
-The paper has been organised as follows: Sec.2 introduces the notations and the relevant preliminaries.
+Paper Organisation: Sec.2 introduces the notations and the relevant preliminaries.
 Sec.
-3 motivates the analysis of the topological dependence of opinion clusters.
+3 motivates the study of the formation of opinion clusters.
 Sec.
-4 presents a framework to relate final opinions of any set of agents in the graph.
-We introduce the LTP agents and present the conditions for opinion clustering in Sec.
-5.
-Finally, we conclude in Sec.
-6 with insights into future directions.
+4, introduces the LTP agents and presents the conditions for opinion clustering.
+Finally, Sec.
+5 concludes with insights into future directions.
 2 Notations and Preliminaries
-2.1 Notations
 […]
-2.2 Graph Preliminaries
 […]
@@
 A cycle is a walk whose initial and final nodes coincide.
-If no nodes in a walk are repeated, it is a path.
+A walk in which no node is repeated is a path.
+A path from node $x$ to node $y$ traverses node $z$ if the sequence of nodes that form this path contains $z$.
 A graph is aperiodic if the lengths of all its cycles are coprime.
@@
 A directed graph (digraph) is a strongly connected graph if a directed path exists between every pair of nodes in the graph.
-A digraph is weakly connected if it is not strongly connected, but its undirected version is connected.
+A digraph is weakly connected if the undirected version of the digraph is connected.
 A maximal subgraph of $\mathcal{G}$ which is strongly connected forms a strongly connected component (SCC).
-An SCC is an independent strongly connected component (iSCC) if every node in the SCC has all its in-neighbours within the same SCC.
-2.3 Kron Reduction
-Let $M\in\mathbb{R}^{p\times p}$ and $\alpha,\beta\subseteq[p]$ be index sets.
-Then $M[\alpha,\beta]$ is a submatrix of $M$ with rows indexed by $\alpha$ and columns indexed by $\beta$.
-The submatrix $M[\alpha]:=M[\alpha,\alpha]$ and $\alpha^{c}:=[p]\setminus\alpha$.
-If $M[\alpha^{c}]$ is non-singular, then the Schur complement of $M[\alpha^{c}]$ in $M$ is defined as:
-$M/\alpha^{c}=M[\alpha]-M[\alpha,\alpha^{c}](M[\alpha^{c}])^{-1}M[\alpha^{c},\alpha]$ (1)
-[…]
-Lemma 1 ([16])
-Consider a loopy Laplacian matrix $Q\in\mathbb{R}^{p}$ with $\alpha\subset[p]$. Let $\mathcal{G}(Q)$ be the digraph associated with $Q$.
-  1. 1.
-    $Q/\alpha^{c}$ is well defined if for each node $i\in\alpha^{c}$ there is a node $j\in\alpha$ such that a path $j\to i$ exists in $\mathcal{G}(Q)$.
-(Note that this condition is adapted to our framework because we consider $w_{ij}>0$ if edge $(j,i)\in\mathcal{E}$.)
-  2. 2.
-    If $Q$ is a loopless Laplacian matrix, then $Q/\alpha^{c}$ is also a loopless Laplacian matrix.
+An SCC is an independent SCC (iSCC) if every node in the SCC has all its in-neighbours within the same SCC.
 3 FJ model and emergent behaviours
 Consider a network $\mathcal{G}$ of $n$ agents with vector $\mathbf{x}(k)=[x_{1}(k),...,x_{n}(k)]$ denoting the opinions of the agents at the $k^{th}$ instance.
-Under the FJ model, the opinions of agents evolve as follows:
+Under the FJ model, the opinions evolve as:
 $\mathbf{x}(k+1)=(I-\beta)W\mathbf{x}(k)+\beta\mathbf{x}(0)$ (1)
@@
 Suppose $\mathcal{G}$ is weakly connected and there is a non-stubborn agent who does not have a directed path from any stubborn agent.
-In [5], such agents are called oblivious agents.
-Without loss of generality, we can renumber the nodes in $\mathcal{G}$ such that the nodes $\{1,2,...,n_{o}\}$ are the oblivious agents and the rest are the non-oblivious, where $n_{o}\in[n]$.
-Additionally, we consider the oblivious agents that form an iSCC to be grouped together.
-Now, we can re-write the FJ-model (2) as:
+Such agents are called oblivious agents [8].
+Without loss of generality, we renumber the nodes in $\mathcal{G}$ such that the nodes $\{1,2,...,n_{o}\}$ are the oblivious agents and the rest are the non-oblivious, where $n_{o}\in[n]$.
+Additionally, we consider the oblivious agents that form an iSCC are grouped together.
+Now, we can re-write the FJ model (1) as:
 $\mathbf{x}_{1}(k+1)$ $=W_{11}\mathbf{x}_{1}(k)$
@@
 where $\mathbf{x}_{1}(k)\in\mathbb{R}^{n_{o}}$ and $\mathbf{x}_{2}(k)\in\mathbb{R}^{(n-n_{o})}$ denote the opinions of oblivious and non-oblivious agents.
-The matrix $W$ gets partitioned as $W=\begin{bmatrix}W_{11}&0\\ W_{21}&W_{22}\end{bmatrix}$ and $\bar{\beta}\in\mathbb{R}^{(n-n_{o})\times(n-n_{o})}$ is a diagonal matrix with entries equal to the stubbornness of the non-oblivious agents.
+The matrix $W$ gets partitioned as $W=\begin{bmatrix}W_{11}&0\\ W_{21}&W_{22}\end{bmatrix}$ and the diagonal matrix $\bar{\beta}$ denotes the stubbornness of non-oblivious agents.
+Throughout this article, we consider that $\mathcal{G}$ satisfies the following:
+Assumption 1
+Each iSCC in $\mathcal{G}$ composed of the oblivious agents with two or more agents is aperiodic.
+Assumption 1 is the necessary and sufficient condition for convergence under the FJ model [9].
 Lemma 1 ([8, 9])
-Consider a weakly connected network $\mathcal{G}$ with opinions evolving under the FJ model (2).
-  • If $\mathcal{G}$ does not contain any oblivious agents, the final opinions $\mathbf{x}^{*}$ always converge to,
-    $\mathbf{x}^{*}=(I-(I-\beta)W)^{-1}\beta\mathbf{x}(0)$
-  • If $\mathcal{G}$ has oblivious agents, then the opinions converge at steady state only if each iSCC composed of oblivious agents is aperiodic. The final opinions converge to,
+Consider a network $\mathcal{G}$ with $m\geq 1$ stubborn agents such that Assumption 1 holds. The opinions evolving under the FJ model (1) converge to,
 $\mathbf{x}_{1}^{*}$ $=W_{11}^{*}\mathbf{x}_{1}(0)$
@@
 $+\bar{\beta}\mathbf{x}_{2}(0)\big)$
-    where $W_{11}^{*}=\lim_{k\to\infty}W_{11}^{k}$.
-Since $\mathbf{x}^{*}$ depends only on the initial opinions of the stubborn and the oblivious agents in iSCCs, these agents are called the influential agents.
-$\mathcal{I}$ denotes the set of all influential agents in $\mathcal{G}$.
-From Lemma 2, it follows that the impact of these influential agents on $\mathbf{x}^{*}$ depends on two factors: the network topology (captured by $W$) and the stubbornness of agents.
-In general, the final opinions converge to different values, causing disagreement among agents.
-However, certain subgroups of agents can still have consensus amongst themselves and form opinion clusters in the final opinion.
-Naturally, the question arises how the formation of these opinion clusters depends on network topology and stubborn behaviour.
-The following example illustrates the effect of network topology on the opinion clusters.
+where $\mathbf{x}^{*}=[\mathbf{x}_{1}^{*},\mathbf{x}_{2}^{*}]$ denotes the final opinions and $W_{11}^{*}=\lim_{k\to\infty}W_{11}^{k}$.
+From Lemma 1, it clearly follows that the final opinions depend on the initial opinions of only the stubborn agents and the oblivious agents in iSCCs.
+Hence, these agents are referred to as the influential agents.
+The set of all influential agents in $\mathcal{G}$ is denoted by $\mathcal{I}$.
+The reference [9] establishes that the final opinions lie in the convex hull of the initial opinions of influential agents.
+In [15], the coefficients of each of these convex combinations (that determine the final opinions) are related to the network topology (captured by $W$) and the stubbornness of agents.
+Depending on these factors, the final opinions of agents (except the oblivious agents in an iSCC) generally converge to distinct values, which is termed as disagreement.
+Disagreement is a more prevalent in social networks as compared to global consensus [6].
+A special case of disagreement occurs when a subgroup of two or more agents converges to the same final opinion and forms an opinion cluster.
+An opinion cluster is defined as:
+Definition 1
+(Opinion Cluster) A set of agents $\mathcal{C}\subset\mathcal{V}$ form an opinion cluster if for any initial conditions $\mathbf{x}(0)\in\mathbb{R}^{n}$
+$x_{i}^{*}-x_{j}^{*}=0\quad\forall\ i,j\in\mathcal{C}$
+where $x_{k}^{*}$ denotes the final opinion of agent $k\in\mathcal{V}$.
+From a sociological perspective, the existence of opinion clusters captures a more realistic distribution of opinions in society, where local consensus occurs in groups (echo chambers) despite widespread disagreement.
+The formation of opinion clusters has been examined in opinion dynamics [3], however, the role of network topology leading in this phenomenon remains less explored.
+In [12], the authors present topology-based conditions for the formation of opinion clusters in digraphs where each cycle has at most one stubborn agent.
+However, these conditions are insufficient to explain the emergence of opinion clusters in the digraph in Example 1 that contains a cycle with multiple stubborn agents.
+Example 1
+Consider a strongly connected network $\mathcal{G}$ in Fig.
+1(a) with stubborn agents $2$ and $6$.
+The cycle in $\mathcal{G}$ formed by agents $1,2$ and $6$ contains two stubborn agents.
+Let the stubbornness of the stubborn agents be $\beta_{2}=0.3$, and $\beta_{6}=0.6$.
+For initial opinions from a uniform distribution on $[0,5]$, the final opinions under the FJ model (1) form three opinion clusters as shown in Fig.
+1(b).
+(a) Network $\mathcal{G}$
+(b) Evolution of opinions $\mathcal{G}$.
+Figure 1: Formation of opinion clusters. Throughout this paper, we represent the stubborn and non-stubborn agents by red and orange nodes, respectively.
+In this work, our objective is to establish conditions under which a subgroup of agents in any arbitrarily connected digraph forms an opinion cluster in the FJ framework.
+Using these conditions, we can use network interactions as a tool to form desired opinion clusters.
+Remark 1
+The agents within an opinion cluster converge to the same final opinions, regardless of initial opinions.
+This implies that each influential agent $s\in\mathcal{I}$ exerts an equal impact on all agents in the cluster.
+Thus, this work also presents the network structures under which a group of agents are equally affected by the influential agents.
+4 Topology-based opinion clustering
+Consider a weakly connected network $\mathcal{G}$ with opinions evolving under the FJ model (1).
+Let $q\in\mathcal{V}\setminus\mathcal{I}$ be a non-influential agent in $\mathcal{G}$.
+The final opinion $x_{q}^{*}$ depends on the initial opinions of only those influential agents that have a path to $q$.
+Let $\mathcal{I}_{q}\subseteq\mathcal{I}$ denote the set of influential agents that have a path to $q$.
+Certain topological properties in a network endow some agents with the special property of being ‘locally’ influential; we refer to them as LTP agents.
+More formally:
+Definition 2
+An agent $p\in\mathcal{V}$ is called an LTP agent if the following hold:
+  1. (i)
+    there exists an agent $q\in\mathcal{V}\setminus\mathcal{I}$ such that every path in $\mathcal{G}$ from each influential agent $s\in\mathcal{I}_{q}$ to $q$ (where $p\neq q$) traverses $p$,
+  2. (ii)
+    additionally, if $p\in\mathcal{V}\setminus\mathcal{I}$, then there does not exist an agent $c\in\mathcal{V}$ such that all paths from each agent $s\in\mathcal{I}_{p}$ to $p$ traverse $c$.
+The non-influential agent $q$ is then said to be persuaded by $p$.
+The set of all non-influential agents persuaded by an LTP agent $p$ is denoted by $\mathcal{N}_{p}$.
+If $p\in\mathcal{I}$ is an influential agent, condition (i) is sufficient to ensure that $p$ is an LTP agent.
+The condition (ii) in Defn.
+2 is applicable only if $p$ is a non-influential agent.
+It conveys that a non-influential agent $p$, persuaded by an LTP agent $c$ (i.e $p\in\mathcal{N}_{c}$), is not an LTP agent itself.
+The following example illustrates which agents are LTP agents and which are the agents it persuades.
 Example 2
@@
 1(a).
-$\mathcal{G}$ has two stubborn agents: $2$ and $6$ with $\beta_{2}=0.3$, and $\beta_{6}=0.6$, and no oblivious agents.
-For initial opinions chosen from a uniform distribution on $[0,10]$, the opinions under the FJ model (2) evolve as shown in Fig.
-1(b).
-The following opinion clusters form in the final opinion: agents $3,4$ and $5$ converge to $7.95$, agents $1,6$ converge to $8.85$ and agent $2$ converges to $6.83$.
-Next, we modify $\mathcal{G}$ by adding edge $(1,5)$ with weight $0.5$ (and reducing the weight of $(4,5)$ to $0.5$ to ensure $W$ remains row-stochastic) as shown in Fig.1(c).
-For the same stubbornness and the initial opinions, Fig.
-1(d) shows that node $5$ forms a new cluster while the rest belong to the same opinion clusters.
-(a) Network $\mathcal{G}$
-(b) Evolution of opinions $\mathcal{G}$.
-(c) Modified network $\mathcal{\hat{G}}$
-(d) Evolution of opinions in $\mathcal{\hat{G}}$
-Figure 1: Simple modification to $\mathcal{G}$ affects the opinion clusters in final opinion.
-Throughout the paper, we represent the stubborn agents and the non-stubborn agents by red nodes and orange nodes, respectively.
-The modified edges in $\mathcal{\hat{G}}$ are highlighted in red.
-Example 1 demonstrates that the underlying topology has a significant impact on formation of opinion clusters in the network.
-Motivated by this, our objective is to establish the topology-based conditions that steer a subgroup of agents in the network to form an opinion cluster in the FJ framework.
-4 The Relation of the final opinions with Network topology
-In this section, we present a framework to relate the final opinions of agents with the underlying network topology.
-Consider a digraph $\mathcal{G}$ consisting of $n$ agents with opinions evolving under the FJ model (2).
-Let the network have $m$ stubborn agents.
-The final opinions satisfy the equations: $\mathbf{x}^{*}=(I-\beta)W\mathbf{x}^{*}+\beta\mathbf{x}(0)$.
-We can rewrite the above-mentioned linear equations as follows:
-$\underbrace{\begin{bmatrix}(I-(I-\beta)W)&-\eta&\\ 0&0&\end{bmatrix}}_{R}\ \underbrace{\begin{bmatrix}\mathbf{x}^{*}\\ \mathbf{x}_{s}(0)\end{bmatrix}}_{z}=\mathbb{0}$ (3)
-where $\mathbf{x}_{s}=[x_{s_{1}}(0),...,x_{s_{m}}(0)]$ and $s_{1},s_{2},...,s_{m}\in[m]$ are the labels of the stubborn agents.
-The matrix $\eta=[\eta_{ij}]\in\mathbb{R}^{n\times m}$ has entry $\eta_{{i}j}=\beta_{i}$ if the stubborn agent $i$ is labelled as $s_{j}$, otherwise $\eta_{ij}=0$ for $i\in[n],j\in[m]$.
-Henceforth, we represent eqn.
-(3) as $Rz=\mathbb{0}$ where $R=[r_{i,j}]\in\mathbb{R}^{(n+m)\times(n+m)}$.
-Remark 1
-The following are the salient properties of $R$: (a) each row sum in $R$ is 0, (b) the diagonal entries are non-negative while the off-diagonal entries are non-positive.
-Thus, $R$ is a Laplacian matrix under the Defn.
-6.3 in [17].
-The digraph $\mathcal{G}(R)$ can be induced from $R$ if an edge $(i,j)$ exists in $\mathcal{G}(R)$ when $r_{ji}<0$.
-The associated graph $\mathcal{G}(R)$ has $n+m$ nodes.
-Being consistent with the indexing of nodes in $\mathcal{G}$, the first $n$ nodes in $\mathcal{G}(R)$ are associated with $x_{i}^{*}$ for $i\in[n]$ and the nodes $n+1$ to $n+m$ are associated with the initial states of the $m$ stubborn agents.
-Note that each node $n+i$ forms a source in $\mathcal{G}(R)$ for $i\in[m]$.
-The source $n+i$ has a single outgoing edge to the node associated with the final opinion of the corresponding stubborn agent.
-For example, in Fig.
-2, the initial opinion of stubborn agent $2$ is associated with node $n+1$ in $\mathcal{G}(R)$, thus, $n+1$ only has the outgoing edge $(n+1,2)$.
-Example 2
-Consider the network $\mathcal{G}$ shown in Fig.
-1(a) with stubborn agents $2$ and $6$.
-The network $\mathcal{G}(R)$ for the matrix $R$ derived from eqn.
-(3) is shown in Fig.
-2.
-The nodes in $\mathcal{G}(R)$ that represent the initial opinion of the stubborn agents are highlighted in turquoise.
-Figure 2: The network $\mathcal{G}(R)$ derived from $\mathcal{G}$.
-Remark 2
-Since $R$ is a Laplacian matrix, it can be reduced by Kron reduction for a suitable choice of the subset of nodes $\alpha$ (defined in Sec.
-2.3).
-Consider a network $\mathcal{G}$ without any oblivious agents.
-For such a network, each node in $[n]$ has a path in $\mathcal{G}(R)$ from at least one source $\{n+1,...,n+m\}$.
-Thus, by Statement 1) in Lemma 1, the Schur complement $R/\alpha^{c}$ is well-defined for any $\alpha$ such that $\{n+1,...,n+m\}\subseteq\alpha\subseteq[n+m]$.
-If $\mathcal{G}$ has an oblivious agent, then the sources $\{n+1,...,n+m\}$ do not have a path to this agent.
-Thus, we define $\alpha$ to include the oblivious agents belonging to each iSCC along with sources $\{n+1,...,n+m\}$.
-For this choice of $\alpha$, each $i\in\alpha^{c}$ has a path in $\mathcal{G}(R)$ from a node $j\in\alpha$.
-Thus, Statement 1) in Lemma 1 holds.
-Importantly, eliminating the states associated with nodes in $\alpha^{c}$ in eqn. (3), reduces it to
-$R/\alpha^{c}.\begin{bmatrix}\mathbf{x}^{*}[\omega]&\mathbf{x}_{s}(0)]\end{bmatrix}^{T}=\mathbb{0}$ (4)
-where $\omega=\alpha\setminus\{n+1,...,n+m\}$.
-Thus, the Kron reduction technique is useful in determining the relations between the final opinions of a subgroup of nodes in $\mathcal{G}$ and quantifying their dependence on the influential agents (stubborn agents and oblivious agents in iSCC).
-Using the Kron-reduction technique, we present topology-based conditions under which the agents in a subgroup have equal final opinions and form an opinion cluster.
-5 Topology-based opinion clustering
-Consider a weakly connected network $\mathcal{G}$ with opinions evolving under the FJ model (2).
-Let $q$ be a non-influential agent in $\mathcal{G}$.
-Then, it is simple to see that the final opinion $x_{q}^{*}$ depends on the initial opinions of only those influential agents that have a path to $q$.
-Let $\mathcal{I}_{q}\subseteq\mathcal{I}$ be the set of all influential agents that have a directed path to $q$.
-Now, we define an LTP agent in terms of paths from agents in $\mathcal{I}_{q}$ to $q$.
-Definition 1
-Consider an agent $p$ in network $\mathcal{G}$.
-If every path in $\mathcal{G}$ from each influential agent $s\in\mathcal{I}_{q}$ to $q$ (where $p\neq q$) traverses $p$, then $p$ is an LTP agent and $q$ is said to be persuaded by $p$.
-Note that an LTP agent itself is not persuaded by any other agent, i.e.
-all paths from agents in $\mathcal{I}_{p}$ to $p$ must not pass through any other agent.
-The set of all non-influential agents persuaded by an LTP agent $p$ is denoted as $\mathcal{N}_{p}$.
-Since $p$ is an LTP agent and $q\in\mathcal{N}_{p}$, it implies that all possible walks from each agent in $\mathcal{I}_{q}$ to $q$ also traverse $p$.
-An LTP agent can be an influential agent (i.e.
-a stubborn agent or an oblivious agent in an iSCC) or a non-influential agent.
-An influential agent is an LTP agent only if each path from the remaining influential agents in $\mathcal{I}_{q}$ to $q$ traverses $p$.
-The following example illustrates the identification of an LTP agent and the agent(s) it persuades.
-Example 3
-Consider the network $\mathcal{G}$ shown in Fig.
-1(a).
-Since $\mathcal{G}$ is strongly connected, only the stubborn agents $2$ and $6$ are influential and have a path to all agents.
-Fig 3(a) demonstrates that each path from $2$ to $4$ (and $5$) in $\mathcal{G}$ traverses through $3$.
-Moreover, the paths from $6$ to $4$ (and $5$) also traverse through $3$ as displayed in Fig.
-3(b).
+Since $\mathcal{G}$ is strongly connected, only the stubborn agents $2$ and $6$ are influential.
+Fig 2(a) demonstrates that each path from $2$ to $4$ (and $5$) in $\mathcal{G}$ traverses $3$.
+Further, the paths from $6$ to $4$ (and $5$) also traverse $3$ as shown in Fig.
+2(b).
 Since all paths from $2$ and $6$ to $3$ do not traverse any common node.
@@
 Similarly, $6$ is also an LTP agent and $\mathcal{N}_{6}=\{1\}$.
+Note that each path from $2$ and $6$ to $5$ traverses $4$. Hence, $4$ satisfies condition 1) in Definition 2 However, since $4\in\mathcal{N}_{3}$, it is not an LTP agent.
 (a) Paths from $2$ to $4$ (and $5$) traverse 33
 (b) Paths from $6$ to $4$ (and $5$) traverse 33
-Figure 3: Node 33 is an LTP agent and it persuades 44 and $5$ as each path from the stubborn agents traverse $3$.
-Suppose a network $\mathcal{G}$ has multiple LTP agents.
-The following question naturally arises: Can an agent be persuaded by two different LTP agents?
-The following Lemma presents an answer to this question.
+Figure 2: Each path from the stubborn agents to 44 and 55 traverses $3$.
+Remark 2
+The notion of an LTP agent is analogous to that of cut vertices in undirected graphs (and strong articulation points in digraphs).
+A node $v$ is a cut vertex (or a strong articulation point in a strongly connected digraph) if and only if there are distinct nodes $x$ and $y$ such that every path from $x$ to $y$ traverses $v$ [19].
+Since LTP agents are defined for digraphs, each LTP agent in a strongly connected digraph is a strong articulation point; however, the converse does not hold (e.g., node $4$ in Example 2).
+The previous discussion describes the properties of the LTP agents.
+In larger networks, however, multiple LTP agents may exist.
+This leads to the question: Can a non-influential agent be persuaded by more than one LTP agent?
 Lemma 2
 Consider a network $\mathcal{G}$ and let $\mathcal{L}$ denote the set of LTP agents in $\mathcal{G}$. Then, $\mathcal{N}_{i}\cap\mathcal{N}_{j}=\emptyset$ for any distinct $i,j\in\mathcal{L}$.
-Proof 5.1.
+Proof 4.1.
 Suppose the contrary is true and there exists a node $q\in\mathcal{N}_{i}\cap\mathcal{N}_{j}$.
@@
 Under C1, every path from an agent in $\mathcal{I}_{q}$ to $j$ also traverses $i$.
+Note that $j$ must be a non-influential agent under C1; otherwise $i$ will not be an LTP agent (as a path $j\to q$ exists that does not traverse $i$).
 Since $j$ has a path to $q$, it follows that $\mathcal{I}_{j}\subseteq\mathcal{I}_{q}$, and hence every path from each node in $\mathcal{I}_{j}$ to $j$ traverses $i$.
-Thus $j\in\mathcal{N}_{i}$, so $j$ cannot be an LTP agent.
+Thus, $j\in\mathcal{N}_{i}$, so $j$ cannot be an LTP agent.
 Equivalently, under C2, we obtain $i\in\mathcal{N}_{j}$, so $i$ is not an LTP agent.
@@
 Similarly, we can show that a path from $s\to q$ exists that does not traverse $j$.
-Since all of the conditions: C1, C2 and C3, lead to a contradiction, it means that the assumption $q\in\mathcal{N}_{i}\cap\mathcal{N}_{j}$ does not hold.
+Since all of the conditions: C1, C2 and C3, lead to a contradiction, the assumption $q\in\mathcal{N}_{i}\cap\mathcal{N}_{j}$ does not hold.
 Lemma 2 establishes that an LTP agent and the agents it persuades form a disjoint group in the network.
-However, there can be certain agents who are neither an LTP agent nor are persuaded by any LTP agent.
+Certain agents in $\mathcal{G}$ can exist who are neither an LTP agent nor are persuaded by any LTP agent.
 Excluding them, we can uniquely partition the remaining agents into disjoint groups, with each group consisting of one LTP agent and its persuaded agents.
-The following results presents the relation between the final opinions of agents within each such group.
+4.1 The role of network topology on opinion formation
+In this subsection, we present a framework that relates the final opinions of agents with the underlying network topology.
+Subsequently, we employ this framework to demonstrate that the topological relation between an LTP agent and the agent(s) it persuades leads to the formation of an opinion cluster.
+Consider a network $\mathcal{G}$ of $n$ agents with opinions evolving under the FJ model (1).
+Let $\mathcal{G}$ have $m$ stubborn agents.
+Then, the final opinions satisfy: $\mathbf{x}^{*}=(I-\beta)W\mathbf{x}^{*}+\beta\mathbf{x}(0)$.
+We can rewrite these linear equations as follows:
+$\underbrace{\begin{bmatrix}(I-(I-\beta)W)&-\eta&\\ 0&0&\end{bmatrix}}_{R}\ \underbrace{\begin{bmatrix}\mathbf{x}^{*}\\ \mathbf{x}_{s}(0)\end{bmatrix}}_{z}=\mathbb{0}$ (2)
+where $\mathbf{x}_{s}=[x_{s_{1}}(0),...,x_{s_{m}}(0)]$ and $s_{1},s_{2},...,s_{m}\in[m]$ are the labels of the $m$ stubborn agents.
+The matrix $\eta=[\eta_{ij}]\in\mathbb{R}^{n\times m}$ has entry $\eta_{{i}j}=\beta_{i}$ if the stubborn agent $i$ is labelled as $s_{j}$, otherwise $\eta_{ij}=0$ for $i\in[n],j\in[m]$.
+Henceforth, we represent eqn.
+(2) as $Rz=\mathbb{0}$ where $R=[r_{i,j}]\in\mathbb{R}^{(n+m)\times(n+m)}$.
+Remark 4.2.
+The following are the salient properties of $R$: (a) each row sum in $R$ is 0, (b) the diagonal entries are non-negative while the off-diagonal entries are non-positive.
+Thus, $R$ is a Laplacian matrix under the Defn.
+6.3 in [20].
+The digraph $\mathcal{G}(R)$ can be induced from $R$ if an edge $(i,j)$ exists in $\mathcal{G}(R)$ when $r_{ji}<0$.
+The associated graph $\mathcal{G}(R)$ has $n+m$ nodes.
+Being consistent with the indexing of nodes in $\mathcal{G}$, the first $n$ nodes in $\mathcal{G}(R)$ are associated with $x_{i}^{*}$ for $i\in[n]$ and the nodes $n+1$ to $n+m$ are associated with the initial states of the $m$ stubborn agents.
+Each node $n+i$ forms a source in $\mathcal{G}(R)$ for $i\in[m]$.
+The source $n+i$ has only one outgoing edge to the node associated with the final opinion of the corresponding stubborn agent.
+For example, in Fig.
+3, the initial opinion of stubborn agent $2$ is associated with node $n+1$ in $\mathcal{G}(R)$, thus, $n+1$ only has outgoing edge $(n+1,2)$.
+Example 3
+Consider the network $\mathcal{G}$ in Fig.
+1(a) with stubborn agents $2$ and $6$.
+The corresponding network $\mathcal{G}(R)$, shown in Fig.
+3, is derived from eqn.
+(2).
+Each node associated with the initial opinion of a stubborn agent is shown in turquoise.
+Figure 3: The network $\mathcal{G}(R)$ derived from $\mathcal{G}$.
+Remark 4.3.
+Under Kron Reduction, we determine the Schur complement (4) of the loopy Laplacian matrix $Q$.
+Since $R$ is a Laplacian matrix, it can be reduced by Kron Reduction for a suitable choice of the subset of nodes $\alpha$.
+If $\mathcal{G}$ does not have any oblivious agents, then each node in $[n]$ has a path in $\mathcal{G}(R)$ from at least one source $\{n+1,...,n+m\}$.
+Thus, by Statement 1) in Lemma 5.1, the Schur complement $R/\alpha^{c}$ is well-defined for any $\alpha$ such that $\{n+1,...,n+m\}\subseteq\alpha\subseteq[n+m]$.
+On the contrary, if $\mathcal{G}$ has an oblivious agent, then the sources $\{n+1,...,n+m\}$ do not have a path in $G(R)$ to this agent.
+For such networks, the set $\alpha$ must include the oblivious agents belonging to each iSCC and sources $\{n+1,...,n+m\}$.
+For this choice of $\alpha$, each $i\in\alpha^{c}$ has a path in $\mathcal{G}(R)$ from a node $j\in\alpha$.
+Thus, Statement 1) in Lemma 5.1 holds.
+Consider a set of linear equations $M\mathbf{y}=\mathbb{0}$, where $\mathbf{y}\in\mathbb{R}^{p}$.
+Eliminating $y[\alpha^{c}]$ using Schur complement (4), yields the reduced system of equations: $M/\alpha^{c}\ \mathbf{y}[\alpha]=0$ (Pg.
+18 in [21]).
+Similarly, we eliminate the states $z[\alpha^{c}]$ in the linear eqns.
+(2) using the Schur Complement to obtain,
+$R/\alpha^{c}.\begin{bmatrix}\mathbf{x}^{*}[\omega]&\mathbf{x}_{s}(0)]\end{bmatrix}^{T}=\mathbb{0}$ (3)
+where $\omega=\alpha\setminus\{n+1,...,n+m\}$.
+Thus, the Kron Reduction technique is useful in determining the relations between the final opinions of a subgroup of nodes in $\mathcal{G}$ and quantifying their dependence on the influential agents.
+4.2 Main Results
+Using the Kron Reduction technique, we present topology-based conditions under which the agents in a subgroup have equal final opinions and form an opinion cluster.
+Theorem 1.
+Consider a network $\mathcal{G}$ with $m\geq 1$ stubborn agents.
+Let Assumption 1 hold.
+The opinions of agents in $\mathcal{G}$ are governed by the FJ model (1).
+If $p$ is an LTP agent, then $p$ and the nodes in set $\mathcal{N}_{p}$ form an opinion cluster.
+The proof is given in the Appendix.
+In [15], the authors show that the final opinion of an agent depends on the paths (the weights of edges on the path) to it from the stubborn agents.
+However, it remains unclear what topological conditions on the paths from stubborn agents to distinct agents ensure that the final opinions of these agents converge to the same value.
+Theorem 1 addresses this gap by proving that the topological relation between an LTP agent and the agents it persuades leads to the formation of an opinion cluster.
+A major advantage of this result is that regardless of edge weights in the network $\mathcal{G}$ and the stubbornness values, an LTP agent $p$ and the agents in $\mathcal{N}_{p}$ collectively form an opinion cluster.
+Example 1 illustrates the result in Theorem 1 as the LTP agent $3$ and agents in $\mathcal{N}_{3}=\{4,5\}$ form an opinion cluster.
+Similarly, the LTP agent $6$ forms another opinion cluster with agent in $\mathcal{N}_{6}=\{1\}$.
+Remark 4.4.
+In [12], the authors derived the conditions to obtain $k\in\mathbb{N}$ opinion clusters in weakly connected digraphs, under the assumption that each cycle in the digraph has at most one stubborn agent.
+Theorem 1 generalises these conditions to form opinion clusters in any arbitrarily connected digraphs, thereby removing this restrictive assumption.
+Moreover, in [12] an opinion cluster is formed by group of agents composed either (1) of a single stubborn agent with the rest being non-stubborn, or (2) of all the oblivious agents in the network $\mathcal{G}$ when $\mathcal{G}$ has a unique aperiodic iSCC of oblivious agents.
+Hence, a group containing only non-influential agents does not form any opinion cluster in [12].
+In contrast, Theorem 1 establishes that an opinion cluster can be formed by any group of non-influential agents if one of them is an LTP agent $p$ and the rest are in $\mathcal{N}_{p}$ (e.g., the agents $3,4,5$ in Example 1).
+Remark 4.5.
+Consider an agent $p\in\mathcal{V}$.
+To verify if $p$ is an LTP agent, we first remove all the outgoing edges of $p$, which takes $O(n)$ time.
+Next, the modified $\mathcal{G}$ is traversed $m+t$ times using the Depth First Search (DFS), once from each of the $m$ stubborn agents and the $t$ iSCCs composed of oblivious agents.
+As one iteration of DFS has complexity $O(n+|\mathcal{E}|)$, this step requires $O((m+t)(n+|\mathcal{E}|))$ time, where $|\mathcal{E}|$ denotes the number of edges.
+If a node remains unvisited after all these traversals, then $p$ satisfies Condition (i) in Defn 2.
+Let $\mathcal{U}_{p}$ be the set of unvisited nodes.
+By repeating this step for all $p\in\mathcal{V}$, we identify all agents that satisfy Condition (i) and store them as $\mathcal{H}$.
+This step has a time complexity of $O((t+m)n(n+|\mathcal{E}|))$ which reduces to $O(n(n+|\mathcal{E}|))$ since $t+m\ll n$.
+If $\mathcal{H}\subseteq\mathcal{I}$, identifying all LTP agents requires $O(n(n+|\mathcal{E}|))$ time.
+Next, we verify condition (ii) in Defn.
+2 for each non-influential agent in $\mathcal{H}$.
+For an non-influential agent $p$, we check if there is a $c\in\mathcal{H}\setminus\{p\}$ such that $p\in\mathcal{U}_{c}$.
+This step requires $O(n^{2})$ time.
+Repeating it for all $p\in\mathcal{H}$ yields a worst-case complexity of $O(n^{3})$.
+Thus, the overall time complexity of identifying all LTP agents in $\mathcal{G}$ is $O(n^{3})$.
+Using the topological conditions in Theorems 1, a network can be designed to obtain a predefined number of opinion clusters with each cluster formed by a desired set of agents.
+Corollary 2.
+Consider a network $\mathcal{G}$ with $m\geq 1$ stubborn agents and $t\geq 0$ iSCCs with oblivious agents such that Assumption 1 holds.
+The opinions of agents in $\mathcal{G}$ are governed by the FJ model (1).
+Let $\mathcal{O}$ be the set of the oblivious agents that either belong to an iSCC or have paths from only one iSCC in $\mathcal{G}$.
+Then, $k\geq t+m$ opinion clusters form in $\mathbf{x}^{*}$ if the agents in set $\mathcal{V}\setminus\mathcal{O}$ can be partitioned into ktk-t disjoint groups with each group composed of an LTP agent and the nodes persuaded by it.
+The proof of Corollary 2 is given in the Appendix. The following example illustrates the conditions presented in Corollary 2 to obtain the desired opinion clustering.
+Example 4
+Consider the network $\mathcal{G}$ with stubborn agents $8$ and $10$ in Fig.
+4(a).
+The agents $1,2$ and $3$ are oblivious agents which together form an iSCC (highlighted by a dashed red circle); the remaining agents are non-oblivious.
+The non-oblivious agents in $\mathcal{G}$ form $4$ disjoint groups highlighted by the green dashed boxes in Fig.
+4(a).
+Each group has an LTP agent (the node with green boundary) and the remaining agents in the group are persuaded by it.
+For initial conditions from a uniform distribution over $[0,10]$ and stubbornness values in $[0,1)$, the opinions of agents in $\mathcal{G}$ under the FJ model form $5$ opinion clusters as shown in Fig.
+4(b).
+One cluster is formed by the oblivious agent belonging to the iSCC and the rest are formed by the agents belonging to the same group.
+(a) Network $\mathcal{G}$
+(b) Formation of desired opinion clusters
+Figure 4: Suitable design of $\mathcal{G}$ for desired opinion clusters.
+5 Conclusions
+The FJ model incorporates the individual biases as stubborn behaviour in the averaging-based opinion models.
+Due to varied biases, even the individuals in a closely connected group hold diverse opinions.
+Even then, a subgroup of individuals can still reach consensus, thereby, forming an opinion cluster.
+This work examines the role of the underlying network topology in the emergence of these clusters.
+To begin with, we define a topologically special agent called an LTP agent, and a set of non-influential agents that it persuades.
+Every path from an influential agent (a stubborn agent or an oblivious agent in an iSCC) to the persuaded agent always traverses the LTP agent.
+Using Kron Reduction, we establish in Theorem 1 that an LTP agent and the set of agents it persuades form an opinion cluster in any arbitrarily-connected digraph under the FJ framework.
+Some additional key insights from our results are as follows: (i) The presence of LTP agents always results in opinion clusters, independent of the edge-weights in the network and the stubbornness of the agents.
+(ii) The notion of LTP agents generalises the topology-based conditions for opinion clustering in [12] to any arbitrary digraphs.
+(iii) By suitably placing the LTP agents and designing the network topology, we achieve any desired opinion clustering.
+In future, we plan to use the LTP agents for desirably shaping the final opinions in real-world social networks.
+Appendix
+Kron Reduction: Let $M\in\mathbb{R}^{p\times p}$ with index set $\alpha\subseteq[p]$ and $\alpha^{c}:=[p]\setminus\alpha$.
+If $M[\alpha^{c}]$ is non-singular, then the Schur complement of $M[\alpha^{c}]$ in $M$ is defined as:
+$M/\alpha^{c}=M[\alpha]-M[\alpha,\alpha^{c}](M[\alpha^{c}])^{-1}M[\alpha^{c},\alpha]$ (4)
+Under Kron Reduction, the Schur complement of the loopy Laplacian matrix $Q$ is evaluated.
+Lemma 5.1 ([18]).
+Consider a loopy Laplacian matrix $Q\in\mathbb{R}^{p}$ with $\alpha\subset[p]$. Let $\mathcal{G}(Q)$ be the digraph associated with $Q$.
+  1. 1.
+    $Q/\alpha^{c}$ is well defined ( i.e.
+$Q[\alpha^{c}]$ is invertible) if for each node $i\in\alpha^{c}$ there is a node $j\in\alpha$ such that a path $j\to i$ exists in $\mathcal{G}(Q)$.
+111(Note that we adapt this condition to our framework because we consider $w_{ij}>0$ when $(j,i)\in\mathcal{E}$.)
+  2. 2.
+    If $Q$ is a loopless Laplacian matrix, then $Q/\alpha^{c}$ is also a loopless Laplacian matrix.
+5.1 Proof of Theorem-1
+Before moving to the proof of Theorem-1, the following Lemma needs to be stated.
 Lemma 5.2.
-Consider a weakly connected network $\mathcal{G}$ with matrix $R$ derived using eqn.
+Consider a weakly connected network $\mathcal{G}$ with $m\geq 1$ stubborn agents and matrix $R$ derived using eqn.
 (2).
@@
 To prove that $\rho(M)<1$, it suffices to show that (I) $M$ is row substochastic, (II) for each $i^{th}$ row in $M$ with row-sum is $1$ there exists a $j^{th}$ row with row-sum less than $1$ such that there is a path from $j$ to $i$ in the associated digraph $\mathcal{G}(M)$.
-Then, by Theorem 6.37 in [18], $\rho(M)<1$.
-Step-I: $M$ can be row-substochastic because: (i) a node $i\in\alpha^{c}$ whose in-neighbour is $p,q$ or an oblivious agent in iSCC satisfies $\sum_{j\in\alpha^{c}}w_{ij}<1$, (ii) each stubborn agent $i$ has the corresponding row-sum at most equal to $(1-\beta_{i})<1$.
-In an arbitrary digraph, the following scenarios may occur:
+Then, by Theorem 4.11 in [20])1, $\rho(M)<1$.
+Step-I: $M$ is row-substochastic if: (i) a node $i\in\alpha^{c}$ whose in-neighbour is $p,q$ or an oblivious agent in iSCC satisfies $\sum_{j\in\alpha^{c}}w_{ij}<1$, (ii) node $i\in\alpha^{c}$ is a stubborn agent and the corresponding row-sum is at most $(1-\beta_{i})<1$.
+In an arbitrary digraph $\mathcal{G}$, the following scenarios may occur:
   • $h\in\alpha^{c}$ has an oblivious agent: Due to weak connectivity of $\mathcal{G}$, an oblivious agent in iSCC exists in $\alpha$ that has a path to $h$, implying that (i) holds.
@@
 Thus, at least one node exists in $\alpha^{c}$ that has either $p$ or $q$ as in-neighbour, implying that (i) holds.
-Thus, there always exists a node in $\alpha^{c}$ that satisfies either (i) or (ii). It implies that $M$ is row substochastic.
+Thus, there always exists a node in $\alpha^{c}$ that satisfies either (i) or (ii). Hence, $M$ is row substochastic.
 Step-II: As discussed in Step-I, each oblivious agent has a path from an oblivious agent in iSCC in $\mathcal{G}$.
@@
 A special case could occur if each path in $\mathcal{G}$ from the oblivious agents in iSCC to an oblivious agent $\tilde{j}\in\alpha^{c}$ either traverses $p$ or $q$.
-In this case, the in-neighbours of $p$ or $q$ in $\alpha^{c}$ who have row-sum in $M$ less than $1$ lie on this path.
+In this case, the in-neighbour(s) of $p$ or $q$ in $\alpha^{c}$ who has row-sum in $M$ less than $1$ lie on this path.
 Hence, every oblivious agent in $\alpha^{c}$ with row sum $1$ has a path in $\mathcal{G}(M)$ from a node with row sum <1<1.
@@
 Thus, statement II also holds and $\rho(M)<1$.
-Theorem 1.
-Consider a network $\mathcal{G}$ of $n$ agents with $m$ stubborn agents and no oblivious agents.
-The opinions of agents in $\mathcal{G}$ are governed by the FJ model (2).
-If $p$ is an LTP agent, then opinions of $p$ and nodes in set $\mathcal{N}_{p}$ converge to the same final opinion and form an opinion cluster.
-Proof 5.4.
-Let $q\in\mathcal{N}_{p}$.
+Proof of Theorem-1: Let $q\in\mathcal{N}_{p}$.
 To prove that the final opinion $x_{p}^{*}=x_{q}^{*}$, we reduce the matrix $R$ in eqn.
 (2) using the Kron Reduction and examine the relation between $x_{p}^{*}$ and $x_{q}^{*}$ in the reduced set of linear equations (3).
-Since $\mathcal{G}$ does not have any oblivious agents, each non-stubborn agent has a path from at least one stubborn agent.
-From the discussion in Remark 2, we know that the Kron Reduction $R/\alpha^{c}$ is well-defined for $\alpha=\{p,q,n+1,...,n+m\}\supset\{n+1,...,n+m\}$.
-Next, we evaluate the Schur complement $R/\alpha^{c}$ given by eqn.
+Let the first $u$ agents in $\mathcal{G}$ be oblivious agents that form the iSCCs.
+The oblivious agents in each iSCC in $\mathcal{G}$ have consensus [9].
+Hence, we consider $q$ is not an oblivious agent in an iSCC; otherwise, $p$ and $q$ belong to an iSCC that is known to achieve consensus.
+Consider $\alpha$ to be $\{p,q,1,...,u,n+1,...,n+m\}$.
+From Remark 4.3, we know that the Kron Reduction $R/\alpha^{c}$ is well-defined.
+Next, we evaluate $R/\alpha^{c}$ given by eqn.
 (4).
-The following structural properties of the associated network $\mathcal{G}(R)$ will be use in determining $R/\alpha^{c}$,
-  1. 1.
-    Each walk from $s\in\mathcal{I}_{q}$ to $q$ in $\mathcal{G}$ traverses $p$ for all $s\in\mathcal{I}_{q}$.
-Thus, by construction of $\mathcal{G}(R)$, each walk that from $n+i$ to $q$ in $\mathcal{G}(R)$ will traverse $p$ for all $i\in[m]$.
-  2. 2.
-    The nodes n+1,,n+mn+1,...,n+m in $\mathcal{G}(R)$ are sources.
-Property 1) implies that $r_{q,n+i}=0$ for all $i\in[m]$ because node $n+i$ cannot be an in-neighbour of $q$ in $\mathcal{G}(R)$.
-Further, property 2) implies that the rows of $R$ indexed $\{n+1,...,n+m\}$ have all entries as zero.
-Thus,
-$R[\alpha]=\begin{bmatrix}r_{p,p}&r_{p,q}&r_{p,n+1}&...&r_{p,n+m}\\ r_{q,p}&r_{q,q}&0&...&0&\\ \mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&...&\mathbb{0}_{m\times 1}\end{bmatrix}$.
+The following structural property of the associated network $\mathcal{G}(R)$ will be used in determining $R/\alpha^{c}$:
+  1. P1:
+    Each walk that exists from $s\in\mathcal{I}_{q}$ to $q$ in $\mathcal{G}$ traverses $p$ for all $s\in\mathcal{I}_{q}$.
+Thus, by construction of $\mathcal{G}(R)$, each walk that exists from $j$ to $q$ in $\mathcal{G}(R)$ will traverse $p$ for all $j\in\{[u],n+1,...,n+m\}$.
+Property P1 implies that $r_{q,j}=0$ for all $j\in\{[u],n+1,...,n+m\}$ because node $j$ cannot be an in-neighbour of $q$ in $\mathcal{G}(R)$.
+Additionally, the nodes n+1,,n+mn+1,...,n+m in $\mathcal{G}(R)$ are sources.
+Thus, the corresponding rows of $R$ have all entries as zero.
+Therefore, $R[\alpha]$ takes the following form:
+$R[\alpha]={\begin{bmatrix}r_{p,p}&r_{p,q}&R[p,[u]]&r_{p,n+1}&...&r_{p,n+m}\\ r_{q,p}&r_{q,q}&\mathbb{0}&0&...&0&\\ \mathbb{0}&\mathbb{0}&R[[u]]&\mathbb{0}&\cdots&\mathbb{0}\\ \mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&...&\mathbb{0}_{m\times 1}\end{bmatrix}}$.
 Next, we consider $Y=R[\alpha,\alpha^{c}](R[\alpha^{c}])^{-1}R[\alpha^{c},\alpha]$. From Lemma 5.2, it follows that
 $Y=R[\alpha,\alpha^{c}]\sum_{k=0}^{\infty}\big((I-\beta[\alpha^{c}])W[\alpha^{c}]\big)^{k}R[\alpha^{c},\alpha]$ (6)
-Let $F:[m+2]\to\alpha$ be a mapping such that $F(1)=p,F(2)=q$ and $F(j+2)=n+j$ for $j\in[m]$.
-This mapping relates the rows (and columns) indices of $Y$ (and $R/\alpha^{c}$) with the agents in set $\alpha$.
+Let $F:[{u}+m+2]\to\alpha$ be a mapping such that $F(1)=p,F(2)=q$, $F(j+2)=j$ for $j\in[u]$, and $F(u+2+j)=n+j$ for $j\in[m]$.
+This mapping relates the row (and column) indices of $Y$ (and $R/\alpha^{c}$) with the agents in set $\alpha$.
 Then, from eqn.
 (6), the entry $y_{ij}$ of $Y$ is non-zero only if there exists a walk in $\mathcal{G}(R)$ from $F(j)$ to $F(i)$ such that each node on this walk (except $F(i)$ and $F(j)$) belongs to $\alpha^{c}$.
-By property 1), $y_{2j}=0$ for all $j\in\{3,...,m+2\}$ because each walk from n+j2n+j-2 to $q$ in $\mathcal{G}(R)$ always passes through $p\in\alpha$.
-Consequently, $R/\alpha^{c}=R[\alpha]-Y$ has the following form:
-$R/\alpha^{c}=\begin{bmatrix}r_{1,1}^{1}&r_{1,2}^{1}&r_{1,3}^{1}&...&r_{1,m+2}^{1}\\ r_{2,1}^{1}&r_{2,2}^{1}&0&...&0\\ \mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&...&\mathbb{0}_{m\times 1}\end{bmatrix}$ where $r_{i,j}^{1}$ gives the entry of matrix $R/\alpha^{c}$ for $i,j\in[m+2]$.
-From Lemma 1, we know that $R/\alpha^{c}$ is also a Laplacian matrix and the row-sum of each of its rows equals 0.
+By property P1), $y_{2j}=0$ for all $j\in{\{3,...,u+m+2\}}$ because each walk from corresponding $F(j)$ to $q$ in $\mathcal{G}(R)$ always passes through $p\in\alpha$.
+Consequently, $R/\alpha^{c}=R[\alpha]-Y$ has the form:
+$\begin{bmatrix}r_{1,1}^{1}&r_{1,2}^{1}&R_{1}&r_{1,u+2}^{1}&...&r_{1,u+m+2}^{1}\\ r_{2,1}^{1}&r_{2,2}^{1}&\mathbb{0}_{u}&0&...&0\\ \mathbb{0}_{u\times 1}&\mathbb{0}_{u\times 1}&R[[u]]&\mathbb{0}_{u\times 1}&\cdots&\mathbb{0}_{u\times 1}\\ \mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&...&\mathbb{0}_{m\times 1}\end{bmatrix}$}
+where $r_{i,j}^{1}$ gives the entry of matrix $R/\alpha^{c}$ for $i,j\in[m+2]$.
+The matrix $R_{1}$ is the submatrix $R/\alpha^{c}[1,[u]]$ highlighting the corresponding non-zero entries.
+From Lemma 5.1, we know that $R/\alpha^{c}$ is also a Laplacian matrix and the row-sum of each of its rows equals 0.
 Since the second row of $R/\alpha^{c}$ has only two non-zero entries, thus, $r_{2,1}^{1}=-r_{2,2}^{1}$.
-Moreover, by eliminating the states $\alpha^{c}$ in the steady state eqns.
-(3), we get the following reduced system:
-$\begin{bmatrix}r_{1,1}^{1}&r_{1,2}^{1}&r_{1,3}^{1}&...&r_{1,m+2}^{1}\\ -r_{2,2}^{1}&r_{2,2}^{1}&0&...&0\\ \mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&...&\mathbb{0}_{m\times 1}\end{bmatrix}\begin{bmatrix}x_{p}^{*}\\ x_{q}^{*}\\ \mathbf{x}_{s}(0)\end{bmatrix}=\mathbb{0}$ (7)
-Simplifying eqn. (7) shows that $x_{p}^{*}=x_{q}^{*}$. Since $q$ is any node in $\mathcal{N}_{p}$, this holds for all nodes in $\mathcal{N}_{p}$.
-Theorem 1 establishes that each LTP agent $p$ and the agents in $\mathcal{N}_{p}$ collectively form an opinion cluster in the final opinion.
-Importantly, the formation of an opinion cluster depends on the existence of an LTP agent in a network, independent of the edge weights and the stubbornness of the stubborn agents.
-The Example 1 illustrates the result in Theorem 1 as the LTP agent $3$ forms an opinion cluster with agents in $\mathcal{N}_{3}=\{4,5\}$ and LTP agent $6$ forms another opinion cluster with agent in $\mathcal{N}_{6}=\{1\}$.
-Additionally, it explains the formation of a new opinion cluster for the modified graph $\mathcal{\hat{G}}$.
-This occurs because the addition of edge $(1,5)$ results in a path from $6$ to $5$ that does not traverse $3$.
-Thus, $5\notin\mathcal{N}_{3}$ and forms a new opinion cluster in Fig.
-1(d).
-Remark 5.5.
-An LTP agent can be a stubborn or a non-stubborn agent.
-However, it is crucial in ensuring that the nodes in its close vicinity in $\mathcal{N}_{p}$ form an opinion cluster.
-Thus, even a non-influential LTP agent impacts the final opinion formation, making it topologically influential.
-The following result extends Theorem 1 to the networks containing the oblivious agents.
-Theorem 2.
-Consider a network $\mathcal{G}$ of $n$ agents with $m$ stubborn agents and at least one oblivious agent.
-The opinions of agents in $\mathcal{G}$ are governed by the FJ model (2).
-If each iSCC in $\mathcal{G}$ composed of the oblivious agents is aperiodic, then final opinions of an LTP agent $p$ and nodes in set $\mathcal{N}_{p}$ form an opinion cluster.
-The proof is along the same lines as Theorem 1 and is omitted for brevity.
-By definition, if an LTP agent $p$ is oblivious, then the agents in $\mathcal{N}_{p}$ are also oblivious.
-Consequently, it follows from Theorem 2 that a set of non-influential oblivious agents can also form an opinion cluster.
-Since the opinion of an oblivious agent evolves according to DeGroot’s model and is not affected by stubborn behaviour, an LTP agent ensures the formation of opinion clusters even in DeGroot’s framework.
-Note that since nodes in $\mathcal{N}_{q}$ are always non-influential, the desired opinion clusters can be formed in DeGroot’s framework only in weakly connected networks.
-Remark 5.6.
-In [14], a network with $m$ stubborn agents and no oblivious agents has $m$ opinion clusters in the final opinion if the agents can be partitioned into subgraphs such that each subgraph contains only one stubborn agent and has a spanning tree with the stubborn agent as its root.
-Additionally, only the stubborn agent in a subgraph can receive information from the other subgraphs.
-An equivalent result exists for networks with oblivious agents as well.
-Note that it is assumed that each cycle in the network contains only one stubborn agent.
-The topological conditions in [14] imply that all the paths to a node in a subgraph from the rest of the stubborn agents traverse the stubborn agent to the subgraph in which the node belongs.
-Thus, this stubborn agent is effectively the LTP agent for the agents within a subgraph that collectively form an opinion cluster.
-Hence, Theorems 1 and 2 extend the results presented in [14] to any weakly connected digraph.
-Using the topological conditions in Theorems 1 and 2, a network can be designed to obtain a predefined number of opinion clusters with a desired subset of agents forming them.
-The following example highlights this property.
-Example 4
-Consider the network $\mathcal{G}$ with stubborn agents $8$ and $10$ in Fig.
-4(a).
-The final opinion is desired to form opinion clusters with each cluster containing the agents within the dashed boxes in Fig.
-4(a).
-To achieve desired clustering, we design $\mathcal{G}$ such that each dashed box (except the iSCC composed of oblivious agents that already form an opinion cluster) has an LTP agent and the remaining agents are persuaded by it.
-In Fig.
-4(a), the nodes with green boundaries denote the LTP agents.
-For initial conditions from a uniform distribution over $[0,10]$ and stubbornness values in $[0,1)$, the opinions of agents in $\mathcal{G}$ under the FJ model form the desired opinion clusters as shown in Fig.
-4(b)
-(a) Network $\mathcal{G}$
-(b) Formation of desired opinion clusters
-Figure 4: Suitable design of $\mathcal{G}$ for desired opinion clusters.
-6 Conclusions
-The FJ model incorporates the individual biases as stubborn behaviour in the averaging-based opinion models.
-Due to varied biases, even the individuals in a closely connected group hold diverse opinions.
-Even then, a subgroup of individuals can still reach consensus, thereby, forming an opinion cluster.
-This work examines the role of the underlying network topology in the emergence of these clusters.
-To begin with, we define a topologically special agent called an LTP agent, and a set of non-influential agents that it persuades.
-Every path from an influential agent (a stubborn agent or an oblivious agent in an iSCC) to the persuaded agent always traverses the LTP agent.
-Using Kron-reduction, we establish in Theorems 1 and 2 that the final opinions of an LTP agent and the set of agents it persuades are equal.
-Thus, they form an opinion cluster in any arbitrarily-connected digraph under the FJ framework.
-Interestingly, such LTP agents lead to the formation of opinion clusters even under the DeGroot’s model (see Theorem 2).
-Some additional key insights from our results are: (i) The presence of LTP agents always results in opinion clusters, independent of the edge-weights in the network and the stubbornness of the agents.
-(ii) Notably, an LTP agent may or may not be influential, but still shapes the final opinions of the agents it persuades.
-(iii) The notion of LTP agents generalises the topology-based conditions for opinion clustering in [14] to any arbitrary digraphs.
-(iv) By suitably placing the LTP agents and designing the network topology, we achieve any desired opinion clustering as demonstrated in Example 4.
-In future, we plan to examine the potential of LTP agents in desirably shaping the final opinions in real-world social networks.
+Hence, by eliminating the states $\alpha^{c}$ in the steady state eqns.
+(2), we get $x_{p}^{*}=x_{q}^{*}$.
+Since $q$ is any node in $\mathcal{N}_{p}$, this holds for all nodes in $\mathcal{N}_{p}$.
+$\blacksquare$
+5.2 Proof of Corollary 2
+We classified the agents in $\mathcal{G}$ into sets $\mathcal{O}$ and $\mathcal{V}\setminus\mathcal{O}$.
+As $\mathcal{V}\setminus\mathcal{O}$ is union of tkt-k disjoint sets $\{i\}\cup\mathcal{N}_{i}$, where each $i\in\mathcal{L}$ is an LTP agent.
+Hence, ktk-t LTP agents are present in set $\mathcal{V}\setminus\mathcal{O}$.
+By Theorem 1, each LTP agent $i$ and the agents in $\mathcal{N}_{i}$ form an opinion cluster.
+Therefore, the agents in set $\mathcal{V}\setminus\mathcal{O}$ form ktk-t clusters.
+The set $\mathcal{V}\setminus\mathcal{O}$ contains the $m$ stubborn agents.
+Since stubborn agents persuaded by an LTP agent, each stubborn agent must be an LTP agent.
+Thus, $k-t\geq m$.
+[…]
+Let $C_{j}$ be the set of oblivious agents in $j^{th}$ iSCC and those oblivious agents only have a path from the $j^{th}$ iSCC.
+The opinions of agents in $C_{j}$ evolve as $\mathbf{y}_{j}(k)=W_{i}^{j}\mathbf{y}_{j}(k)$ with $\mathbf{y}_{j}$ is the vector that denotes the opinion of agents in this set.
+Since the associated digraph $G(W_{i}^{t})$ is a subgraph of $\mathcal{G}$ containing a unique aperiodic iSCC, the opinions of the agents converge to consensus (Theorem 5.1 in [20])1.
+Hence, the $t$ subgroups, each composed of an iSCC of oblivious agents and those having a path only from the iSCC, form $t$ opinion clusters.
+Collectively, the oblivious agents in $\mathcal{O}$ and the agents in $\mathcal{V}\setminus\mathcal{O}$ form $k$ opinion clusters such that $k\geq t+m$.

### v3

@@
 A key advantage of the proposed result is that the resulting opinion clusters are independent of the edge weights and the stubbornness of the agents.
-Finally, we demonstrate that by suitably placing LTP agents, one can design networks that achieve any desired opinion clustering.
-{IEEEkeywords}
-Opinion dynamics; Friedkin-Johnsen model; Clustering of opinions.
-1 INTRODUCTION
-The recent evolution of social media networks has generated a keen interest among researchers to study the impact of social interactions on public opinion.
-Several mathematical models such as the DeGroot’s model [1], the FJ model [2], the Hegselmann-Krause model [3], and the Biased assimilation model [4], explain complex aspects of human interactions such as individual biases and homophily.
-Among these, the FJ model is most analytically tractable and explains the frequently observed phenomenon of disagreement among closely interacting individuals [5].
-Moreover, it is widely popular because it is empirically validated on real-world datasets [6], [7].
-The opinions evolving under the FJ model converge to disagreement due to the presence of biased individuals, who are called stubborn agents.
-In [2], the authors show that if each non-stubborn agent in the network is reachable from a stubborn agent, the final opinions depend on the initial opinions of only the stubborn agents.
-Reference [8] introduces the notion of oblivious agents, who are non-stubborn and do not have a path from any stubborn agent.
-The algebraic and graph-theoretic conditions that ensure convergence in networks with oblivious agents are presented in [8] and [9], respectively.
-In opinion dynamics, emergent behaviours such as consensus, polarisation and opinion clustering are of interest to explain and achieve a desired social outcome.
-Several works [6, 10, 11, 12] have examined the FJ framework to reach these behaviours.
-In [6], the authors design stubbornness values that lead to behaviours such as consensus and polarisation.
-On the other hand, reference [10] presents topology-based conditions for graphs with two stubborn agents that result in polarisation.
-Furthermore, the authors in [11] examine the FJ model with randomised interactions and reveal that the opinions of non-stubborn agents can exhibit oscillations in presence of stubborn agents.
-In contrast, the present work focuses on opinion clustering, which is less-explored in the FJ framework.
-Opinion clustering occurs when a group of two or more agents converge to the same final opinion.
-It represents swarm behaviour and has applications in task distribution [13] and formation control [14].
-In the sociological framework, opinion clustering represents the formation of groups with aligned views.
-In this work, we examine the network topology towards achieving opinion clustering.
-Recent works [10, 11] in the FJ framework highlight the impact of topology in shaping emergent behaviours.
-More recently, the study of network topology is receiving attention to analyse the effects of recommender algorithms on opinion formation.
-The seminal works [15] and [16] contributed significantly by quantifying the influence of a stubborn agent using hitting probabilities of random walks over the network.
-These studies enabled the design of topology-based interventions (recommendations) for desirably shaping opinions [16, 17].
-In the present work, we further explore the relation between the network topology and final opinions and address the following problem: Under what topology-based conditions, does a set of agents converge to the same final opinion (form an opinion cluster)?
-We leverage this relationship to obtain desired opinion clustering.
+Finally, we demonstrate using simulation results that, by suitably placing LTP agents, one can design networks that achieve any desired opinion clustering.
+I INTRODUCTION
+In recent decades, social media networks have become a major platform for advertising, conducting socio-political campaigns, and even spreading misinformation.
+This evolution has resulted in an interest towards understanding the impact of social interactions on the formation of opinion.
+The popular opinion dynamics, such as DeGroot model [1], the FJ model [2], the Hegselmann-Krause model [3], and the Biased assimilation model [4], capture the complex aspects of human interactions.
+Among these, the FJ model is popular because it captures a diverse range of emergent behaviours.
+Its popularity also stems from its analytical tractability and performance over large datasets [5].
+The FJ model was proposed in [2] to justify the disagreement among closely interacting individuals.
+The opinions evolving under the FJ model converge to disagreement due to the presence of biased individuals, who are commonly known as the stubborn agents.
+In [2], the authors show that if each non-stubborn agent in the network has a directed path from a stubborn agent, the opinions converge asymptotically at a steady state.
+Further, they demonstrate that the final opinions depend on the initial opinions of only the stubborn agents and lie in the convex hull of these opinions.
+The authors in [6] introduced the notion of oblivious agents, who are non-stubborn and do not have a path from any stubborn agent.
+The algebraic and graph-theoretic conditions that ensure convergence in networks having oblivious agents were presented in [6] and [7], respectively.
+Additionally, in [7], the authors show that the final opinions depend on the initial opinions of both the stubborn agents and some oblivious agents.
+In opinion dynamics, emergent behaviours such as consensus and opinion clustering are of keen interest to explain and achieve a desired social outcome.
+While consensus is being widely explored even in the FJ framework [8], this work focuses on opinion clustering.
+A network achieves opinion clustering when there is a group of two or more agents such that, for any initial opinions, the final opinions of agents in the group are equal.
+Opinion clustering represents swarm behaviour [9] and has applications in task distribution [10] and formation control [11].
+The works [12, 13, 14] present graph-theoretic conditions under which opinion clustering occurs in continuous-time linear invariant systems.
+While it is known that disagreement occurs in the FJ framework, the conditions for opinion clustering remain less explored.
 In [15], the authors present the graph-theoretic conditions to achieve opinion clustering in the FJ framework in networks with oblivious agents.
-However, the result applies only to networks where each cycle contains at most one stubborn agent.
+However, the results apply only to networks where each cycle contains only one stubborn agent.
 Contrary to [15], our work aims to explain the emergence of opinion clusters in the FJ framework in arbitrarily connected digraphs.
-With this, we now highlight our major contributions:
-  • We define the notion of the Locally Topologically Persuasive (LTP) agent. Each LTP agent is associated with a unique set of non-influential agents in its vicinity.
-  • We show that an LTP agent and the agents associated with it always form an opinion cluster under the FJ model.
+In opinion dynamics, an opinion cluster at steady state represents a group of individuals with a shared view on a topic.
+Our focus is specifically on how the network topology directs a certain group of agents to form an opinion cluster.
+With this, we now highlight our major contributions as follows:
+  • We define a new type of agent called the Locally Topologically Persuasive (LTP) agent. We show that each LTP agent is associated with a unique set of non-influential agents in its vicinity.
+  • We establish that an LTP agent and the agents associated with it always form an opinion cluster under the FJ model.
 This result generalises the topology-based conditions presented in [15] to arbitrarily connected digraphs.
-Paper Organisation: Sec.2 introduces the notations and the relevant preliminaries.
+  • Unlike [15], the proposed LTP-based conditions for opinion clustering show that even a group of non-influential agents can form an opinion cluster
+The organisation of the paper is as follows: Sec.
+II introduces the notations and the relevant preliminaries.
 Sec.
-3 motivates the study of the formation of opinion clusters.
+III motivates the analysis of the topological dependence of opinion clusters.
 Sec.
-4, introduces the LTP agents and presents the conditions for opinion clustering.
-Finally, Sec.
-5 concludes with insights into future directions.
-2 Notations and Preliminaries
+IV presents a framework to relate final opinions of any set of agents in the graph.
+We introduce the LTP agents and present the conditions for opinion clustering in Sec.
+V.
+Finally, we conclude in Sec.
+VI.
+II Notations and Preliminaries
+II-A Notations
 […]
+II-B Graph Preliminaries
 […]
@@
 A cycle is a walk whose initial and final nodes coincide.
-A walk in which no node is repeated is a path.
-A path from node $x$ to node $y$ traverses node $z$ if the sequence of nodes that form this path contains $z$.
+If no nodes in a walk are repeated, it is a path.
 A graph is aperiodic if the lengths of all its cycles are coprime.
@@
 A directed graph (digraph) is a strongly connected graph if a directed path exists between every pair of nodes in the graph.
-A digraph is weakly connected if the undirected version of the digraph is connected.
+A digraph is weakly connected if it is not strongly connected, but its undirected version is connected.
 A maximal subgraph of $\mathcal{G}$ which is strongly connected forms a strongly connected component (SCC).
-An SCC is an independent SCC (iSCC) if every node in the SCC has all its in-neighbours within the same SCC.
-3 FJ model and emergent behaviours
+An SCC is an independent strongly connected component (iSCC) if every node in the SCC has all its in-neighbours within the same SCC.
+II-C Kron reduction
+Let $M\in\mathbb{R}^{p\times p}$ and $\alpha,\beta\subseteq[p]$ be index sets.
+Then $M[\alpha,\beta]$ is a submatrix of $M$ with rows indexed by $\alpha$ and columns indexed by $\beta$.
+The submatrix $M[\alpha]:=M[\alpha,\alpha]$ and $\alpha^{c}:=[p]\setminus\alpha$.
+If $M[\alpha^{c}]$ is non-singular, then the Schur complement of $M[\alpha^{c}]$ in $M$ is defined as:
+$M/\alpha^{c}=M[\alpha]-M[\alpha,\alpha^{c}](M[\alpha^{c}])^{-1}M[\alpha^{c},\alpha]$ (1)
+Under Kron reduction, the Schur complement of the loopy Laplacian matrix $Q$ is determined.
+Lemma 1 ([16])
+Consider a loopy Laplacian matrix $Q\in\mathbb{R}^{p}$ and $\mathcal{G}(Q)$ be the digraph associated with $Q$. Let $\alpha\subset[p]$.
+  1. 1.
+    $Q/\alpha^{c}$ is well defined if, for each node $i\in\alpha^{c}$, there is a node $j\in\alpha$ such that a path $j\to i$ exists in $\mathcal{G}(Q)$.
+111(Note that this condition is adapted to our framework because we consider $w_{ij}>0$ when $(j,i)\in\mathcal{E}$.)
+  2. 2.
+    If $Q$ is a loopless Laplacian matrix, then $Q/\alpha^{c}$ is also a loopless Laplacian matrix.
+III FJ model and emergent behaviours
 Consider a network $\mathcal{G}$ of $n$ agents with vector $\mathbf{x}(k)=[x_{1}(k),...,x_{n}(k)]$ denoting the opinions of the agents at the $k^{th}$ instance.
-Under the FJ model, the opinions evolve as:
+Under the FJ model, the opinions of agents evolve as follows:
 $\mathbf{x}(k+1)=(I-\beta)W\mathbf{x}(k)+\beta\mathbf{x}(0)$ (2)
 […]
-Suppose $\mathcal{G}$ is weakly connected and there is a non-stubborn agent who does not have a directed path from any stubborn agent.
-Such agents are called oblivious agents [8].
-Without loss of generality, we renumber the nodes in $\mathcal{G}$ such that the nodes $\{1,2,...,n_{o}\}$ are the oblivious agents and the rest are the non-oblivious, where $n_{o}\in[n]$.
-Additionally, we consider the oblivious agents that form an iSCC are grouped together.
-Now, we can re-write the FJ model (1) as:
+Suppose $\mathcal{G}$ is weakly connected and there is a non-stubborn agent that does not have a directed path from any stubborn agent.
+In [6], such agents are called oblivious agents.
+Without loss of generality, we can renumber the nodes in $\mathcal{G}$ such that the nodes $\{1,2,...,n_{o}\}$ are the oblivious agents and the rest are the non-oblivious, where $n_{o}\in[n]$.
+Additionally, we consider the oblivious agents that form an iSCC to be grouped together.
+Now, we can re-write the FJ-model (2) as:
 $\mathbf{x}_{1}(k+1)$ $=W_{11}\mathbf{x}_{1}(k)$
@@
 where $\mathbf{x}_{1}(k)\in\mathbb{R}^{n_{o}}$ and $\mathbf{x}_{2}(k)\in\mathbb{R}^{(n-n_{o})}$ denote the opinions of oblivious and non-oblivious agents.
-The matrix $W$ gets partitioned as $W=\begin{bmatrix}W_{11}&0\\ W_{21}&W_{22}\end{bmatrix}$ and the diagonal matrix $\bar{\beta}$ denotes the stubbornness of non-oblivious agents.
-Throughout this article, we consider that $\mathcal{G}$ satisfies the following:
-Assumption 1
-Each iSCC in $\mathcal{G}$ composed of the oblivious agents with two or more agents is aperiodic.
-Assumption 1 is the necessary and sufficient condition for convergence under the FJ model [9].
+The matrix $W$ gets partitioned as $W=\begin{bmatrix}W_{11}&\mathbb{0}\\ W_{21}&W_{22}\end{bmatrix}$ and $\bar{\beta}\in\mathbb{R}^{(n-n_{o})\times(n-n_{o})}$ is a diagonal matrix with entries equal to the stubbornness of the non-oblivious agents.
 Lemma 2 ([6, 7])
-Consider a network $\mathcal{G}$ with $m\geq 1$ stubborn agents such that Assumption 1 holds. The opinions evolving under the FJ model (1) converge to,
+Consider a weakly connected network $\mathcal{G}$ with opinions evolving under the FJ model (2).
+  • If $\mathcal{G}$ does not contain any oblivious agents, the final opinions $\mathbf{x}^{*}$ always converge to
+    $\mathbf{x}^{*}=(I-(I-\beta)W)^{-1}\beta\mathbf{x}(0)$
+  • If $\mathcal{G}$ has oblivious agents, then the opinions converge at steady state only if each iSCC composed of two or more oblivious agents is aperiodic.
+The final opinions $\mathbf{x}^{*}$ converge to
     $\mathbf{x}_{1}^{*}$ $=W_{11}^{*}\mathbf{x}_{1}(0)$
@@
     $+\bar{\beta}\mathbf{x}_{2}(0)\big)$
-where $\mathbf{x}^{*}=[\mathbf{x}_{1}^{*},\mathbf{x}_{2}^{*}]$ denotes the final opinions and $W_{11}^{*}=\lim_{k\to\infty}W_{11}^{k}$.
-From Lemma 1, it clearly follows that the final opinions depend on the initial opinions of only the stubborn agents and the oblivious agents in iSCCs.
-Hence, these agents are referred to as the influential agents.
-The set of all influential agents in $\mathcal{G}$ is denoted by $\mathcal{I}$.
-The reference [9] establishes that the final opinions lie in the convex hull of the initial opinions of influential agents.
-In [15], the coefficients of each of these convex combinations (that determine the final opinions) are related to the network topology (captured by $W$) and the stubbornness of agents.
-Depending on these factors, the final opinions of agents (except the oblivious agents in an iSCC) generally converge to distinct values, which is termed as disagreement.
-Disagreement is a more prevalent in social networks as compared to global consensus [6].
-A special case of disagreement occurs when a subgroup of two or more agents converges to the same final opinion and forms an opinion cluster.
-An opinion cluster is defined as:
+    where $W_{11}^{*}=\lim_{k\to\infty}W_{11}^{k}$.
+Since $\mathbf{x}^{*}$ depends only on the initial opinions of the stubborn and the oblivious agents in iSCCs, these agents are called the influential agents.
+$\mathcal{I}$ denotes the set of all influential agents in $\mathcal{G}$.
+From Lemma 2, it follows that the impact of these influential agents on $\mathbf{x}^{*}$ depends on two factors: the network topology (captured by $W$) and the stubbornness of agents.
+In general, the final opinions converge to different values, causing disagreement among agents.
+A special case of disagreement occurs when a group of two or more agents converges to the same final opinion and forms an opinion cluster.
+Mathematically, an opinion cluster is defined as:
 Definition 1
-(Opinion Cluster) A set of agents $\mathcal{C}\subset\mathcal{V}$ form an opinion cluster if for any initial conditions $\mathbf{x}(0)\in\mathbb{R}^{n}$
+A set of agents $\mathcal{C}\subset\mathcal{V}$ form an opinion cluster if for any initial conditions $\mathbf{x}(0)\in\mathbb{R}^{n}$
 $x_{i}^{*}-x_{j}^{*}=0\quad\forall\ i,j\in\mathcal{C}$
 where $x_{k}^{*}$ denotes the final opinion of agent $k\in\mathcal{V}$.
-From a sociological perspective, the existence of opinion clusters captures a more realistic distribution of opinions in society, where local consensus occurs in groups (echo chambers) despite widespread disagreement.
-The formation of opinion clusters has been examined in opinion dynamics [3], however, the role of network topology leading in this phenomenon remains less explored.
-In [12], the authors present topology-based conditions for the formation of opinion clusters in digraphs where each cycle has at most one stubborn agent.
-However, these conditions are insufficient to explain the emergence of opinion clusters in the digraph in Example 1 that contains a cycle with multiple stubborn agents.
+Naturally, the question arises as to how the formation of these opinion clusters depends on network topology and stubborn behaviour.
+The following example illustrates the effect of network topology on the opinion clusters.
 Example 1
-Consider a strongly connected network $\mathcal{G}$ in Fig.
-1(a) with stubborn agents $2$ and $6$.
-The cycle in $\mathcal{G}$ formed by agents $1,2$ and $6$ contains two stubborn agents.
-Let the stubbornness of the stubborn agents be $\beta_{2}=0.3$, and $\beta_{6}=0.6$.
-For initial opinions from a uniform distribution on $[0,5]$, the final opinions under the FJ model (1) form three opinion clusters as shown in Fig.
+Consider the network $\mathcal{G}$ shown in Fig.
+1(a).
+$\mathcal{G}$ has two stubborn agents: $2$ and $6$ with $\beta_{2}=0.3$, and $\beta_{6}=0.6$, and no oblivious agents.
+For initial opinions chosen from a uniform distribution on $[0,10]$, three opinion clusters form under the FJ model (2) as shown in Fig.
 1(b).
+The agents $3,4$ and $5$ form an opinion cluster, agents $1$ and $6$ form another cluster and the agent $2$ forms the third cluster.
+Next, we modify $\mathcal{G}$ by adding edge $(1,5)$ with weight $0.5$ (and reducing the weight of $(4,5)$ to $0.5$ to ensure $W$ remains row-stochastic) as shown in Fig.1(c).
+For the same stubbornness and the initial opinions, Fig.
+1(d) shows that node $5$ forms a new cluster while the rest belong to the same opinion clusters.
 (a) Network $\mathcal{G}$
 (b) Evolution of opinions $\mathcal{G}$.
-Figure 1: Formation of opinion clusters. Throughout this paper, we represent the stubborn and non-stubborn agents by red and orange nodes, respectively.
-In this work, our objective is to establish conditions under which a subgroup of agents in any arbitrarily connected digraph forms an opinion cluster in the FJ framework.
-Using these conditions, we can use network interactions as a tool to form desired opinion clusters.
+(c) Modified network $\mathcal{\hat{G}}$
+(d) Evolution of opinions in $\mathcal{\hat{G}}$
+Figure 1: Simple modification to $\mathcal{G}$ affects the opinion clusters in final opinion.
+Throughout the paper, we represent the stubborn agents and the non-stubborn agents by red nodes and orange nodes, respectively.
+The modified edges in $\mathcal{\hat{G}}$ are highlighted in red.
+Example 1 demonstrates that the underlying topology has a significant impact on formation of opinion clusters in the network.
+Motivated by this, our objective is to establish the topology-based conditions that steer a subgroup of agents in the network to form an opinion cluster in the FJ framework.
+IV The Relation of the final opinions with Network topology
+In this section, we present a framework to relate the final opinions of agents with the underlying network topology.
+Consider a digraph $\mathcal{G}$ consisting of $n$ agents with opinions evolving under the FJ model (2).
+Let the network have $m$ stubborn agents.
+The final opinions satisfy the equations: $\mathbf{x}^{*}=(I-\beta)W\mathbf{x}^{*}+\beta\mathbf{x}(0)$.
+We can rewrite the above-mentioned linear equations as follows:
+$\underbrace{\begin{bmatrix}(I-(I-\beta)W)&-\eta&\\ \mathbb{0}&\mathbb{0}&\end{bmatrix}}_{R}\ \underbrace{\begin{bmatrix}\mathbf{x}^{*}\\ \mathbf{x}_{s}(0)\end{bmatrix}}_{z}=\mathbb{0}$ (3)
+where $\mathbf{x}_{s}=[x_{s_{1}}(0),...,x_{s_{m}}(0)]$ and $s_{1},s_{2},...,s_{m}\in[m]$ are the labels of the stubborn agents.
+The matrix $\eta=[\eta_{ij}]\in\mathbb{R}^{n\times m}$ has entry $\eta_{{i}j}=\beta_{i}$ if the stubborn agent $i$ is labelled as $s_{j}$, otherwise $\eta_{ij}=0$ for $i\in[n],j\in[m]$.
+Henceforth, we represent eqn.
+(3) as $Rz=\mathbb{0}$ where $R=[r_{i,j}]\in\mathbb{R}^{(n+m)\times(n+m)}$.
 Remark 1
-The agents within an opinion cluster converge to the same final opinions, regardless of initial opinions.
-This implies that each influential agent $s\in\mathcal{I}$ exerts an equal impact on all agents in the cluster.
-Thus, this work also presents the network structures under which a group of agents are equally affected by the influential agents.
-4 Topology-based opinion clustering
+The following are the salient properties of $R$: (a) each row sum in $R$ is 0, (b) the diagonal entries are non-negative while the off-diagonal entries are non-positive.
+Thus, $R$ is a Laplacian matrix under the Defn.
+6.3 in [17].
+The digraph $\mathcal{G}(R)$ can be induced from $R$ if an edge $(i,j)$ exists in $\mathcal{G}(R)$ when $r_{ji}<0$.
+The associated graph $\mathcal{G}(R)$ has $n+m$ nodes.
+Being consistent with the indexing of nodes in $\mathcal{G}$, the first $n$ nodes in $\mathcal{G}(R)$ are associated with $x_{i}^{*}$ for $i\in[n]$ and the nodes $n+1$ to $n+m$ are associated with the initial states of the $m$ stubborn agents.
+Note that each node $n+i$ forms a source in $\mathcal{G}(R)$ for $i\in[m]$.
+The source $n+i$ has a single outgoing edge to the node associated with the final opinion of the corresponding stubborn agent.
+For example, in Fig.
+2, the initial opinion of stubborn agent $2$ is associated with node $n+1$ in $\mathcal{G}(R)$, thus, $n+1$ only has the outgoing edge $(n+1,2)$.
+Example 2
+Consider the network $\mathcal{G}$ shown in Fig.
+1(a) with stubborn agents $2$ and $6$.
+The network $\mathcal{G}(R)$ for the matrix $R$ derived from eqn.
+(3) is shown in Fig.
+2.
+The nodes in $\mathcal{G}(R)$ that represent the initial opinion of the stubborn agents are highlighted in turquoise.
+Figure 2: The network $\mathcal{G}(R)$ derived from $\mathcal{G}$.
+Remark 2
+Since $R$ is a Laplacian matrix, it can be reduced by Kron reduction for a suitable choice of the subset of nodes $\alpha$ (defined in Sec.
+II-C).
+Consider a network $\mathcal{G}$ without any oblivious agents.
+For such a network, each node in $[n]$ has a path in $\mathcal{G}(R)$ from at least one source $\{n+1,...,n+m\}$.
+Thus, by Statement 1) in Lemma 1, the Schur complement $R/\alpha^{c}$ is well-defined for any $\alpha$ such that $\{n+1,...,n+m\}\subseteq\alpha\subseteq[n+m]$.
+If $\mathcal{G}$ has an oblivious agent, then the sources $\{n+1,...,n+m\}$ do not have a path to this agent.
+Thus, we define $\alpha$ to include the oblivious agents belonging to each iSCC along with sources $\{n+1,...,n+m\}$.
+For this choice of $\alpha$, each $i\in\alpha^{c}$ has a path in $\mathcal{G}(R)$ from a node $j\in\alpha$.
+Thus, Statement 1) in Lemma 1 holds.
+Importantly, eliminating the states associated with nodes in $\alpha^{c}$ in eqn. (3), reduces it to
+$R/\alpha^{c}.\begin{bmatrix}\mathbf{x}^{*}[\omega]&\mathbf{x}_{s}(0)]\end{bmatrix}^{T}=\mathbb{0}$ (4)
+where $\omega=\alpha\setminus\{n+1,...,n+m\}$.
+Thus, the Kron reduction technique is useful in determining the relations between the final opinions of a subgroup of nodes in $\mathcal{G}$ and quantifying their dependence on the influential agents (stubborn agents and oblivious agents in iSCC).
+Using the Kron reduction technique, we present topology-based conditions under which the agents in a subgroup have equal final opinions and form an opinion cluster.
+V Topology-based opinion clustering
 Consider a weakly connected network $\mathcal{G}$ with opinions evolving under the FJ model (2).
-Let $q\in\mathcal{V}\setminus\mathcal{I}$ be a non-influential agent in $\mathcal{G}$.
-The final opinion $x_{q}^{*}$ depends on the initial opinions of only those influential agents that have a path to $q$.
-Let $\mathcal{I}_{q}\subseteq\mathcal{I}$ denote the set of influential agents that have a path to $q$.
-Certain topological properties in a network endow some agents with the special property of being ‘locally’ influential; we refer to them as LTP agents.
-More formally:
+Let $q$ be a non-influential agent in $\mathcal{G}$.
+Then, it is simple to see that the final opinion $x_{q}^{*}$ depends on the initial opinions of influential agents that have a path to $q$.
+Let $\mathcal{I}_{q}\subseteq\mathcal{I}$ be the set of all influential agents that have a directed path to $q$.
+Now, we define an LTP agent in terms of paths from agents in $\mathcal{I}_{q}$ to $q$.
 Definition 2
-An agent $p\in\mathcal{V}$ is called an LTP agent if the following hold:
+An agent $p\in\mathcal{V}$ is called an LTP agent if the following holds:
   1. (i)
@@
   2. (ii)
-    additionally, if $p\in\mathcal{V}\setminus\mathcal{I}$, then there does not exist an agent $c\in\mathcal{V}$ such that all paths from each agent $s\in\mathcal{I}_{p}$ to $p$ traverse $c$.
-The non-influential agent $q$ is then said to be persuaded by $p$.
-The set of all non-influential agents persuaded by an LTP agent $p$ is denoted by $\mathcal{N}_{p}$.
-If $p\in\mathcal{I}$ is an influential agent, condition (i) is sufficient to ensure that $p$ is an LTP agent.
+    additionally, if $p\in\mathcal{V}\setminus\mathcal{I}$, an agent $c\in\mathcal{V}$ such that all paths from each agent $s\in\mathcal{I}_{p}$ to $p$ traverse $c$ must not exist.
+For an LTP agent $p$, the non-influential agent $q$ satisfying condition (i) is said to be persuaded by $p$.
+The set of all non-influential agents persuaded by an LTP agent $p$ is denoted as $\mathcal{N}_{p}$.
+Note that if $p\in\mathcal{I}$ is an influential agent, condition (i) alone ensures that $p$ is an LTP agent.
 The condition (ii) in Defn.
-2 is applicable only if $p$ is a non-influential agent.
-It conveys that a non-influential agent $p$, persuaded by an LTP agent $c$ (i.e $p\in\mathcal{N}_{c}$), is not an LTP agent itself.
-The following example illustrates which agents are LTP agents and which are the agents it persuades.
+2 is required to be satisfied only when $p$ is non-influential.
+It implies that if a non-influential agent $p$ is persuaded by an LTP agent $c$ (i.e $p\in\mathcal{N}_{c}$), then $p$ itself cannot be an LTP agent.
+The following example illustrates the identification of an LTP agent and the agent(s) it persuades.
 Example 3
@@
 1(a).
-Since $\mathcal{G}$ is strongly connected, only the stubborn agents $2$ and $6$ are influential.
-Fig 2(a) demonstrates that each path from $2$ to $4$ (and $5$) in $\mathcal{G}$ traverses $3$.
-Further, the paths from $6$ to $4$ (and $5$) also traverse $3$ as shown in Fig.
-2(b).
-Since all paths from $2$ and $6$ to $3$ do not traverse any common node.
+Since $\mathcal{G}$ is strongly connected, only the stubborn agents $2$ and $6$ are influential and have a path to all agents.
+Fig 3(a) demonstrates that each path from $2$ to $4$ (and $5$) in $\mathcal{G}$ traverses through $3$.
+Moreover, the paths from $6$ to $4$ (and $5$) also traverse through $3$ as displayed in Fig.
+3(b).
+Since the paths from $2$ and $6$ to $3$ do not traverse any common node.
 Thus, $3$ is an LTP agent and $\mathcal{N}_{3}=\{4,5\}$.
 Similarly, $6$ is also an LTP agent and $\mathcal{N}_{6}=\{1\}$.
-Note that each path from $2$ and $6$ to $5$ traverses $4$. Hence, $4$ satisfies condition 1) in Definition 2 However, since $4\in\mathcal{N}_{3}$, it is not an LTP agent.
+Note that each path from $2$ and $6$ to $5$ traverses $4$. Hence, $4$ satisfies condition 1) in Defn. 2. However, since $4\in\mathcal{N}_{3}$, it is not an LTP agent.
 (a) Paths from $2$ to $4$ (and $5$) traverse 33
 (b) Paths from $6$ to $4$ (and $5$) traverse 33
-Figure 2: Each path from the stubborn agents to 44 and 55 traverses $3$.
-Remark 2
-The notion of an LTP agent is analogous to that of cut vertices in undirected graphs (and strong articulation points in digraphs).
-A node $v$ is a cut vertex (or a strong articulation point in a strongly connected digraph) if and only if there are distinct nodes $x$ and $y$ such that every path from $x$ to $y$ traverses $v$ [19].
-Since LTP agents are defined for digraphs, each LTP agent in a strongly connected digraph is a strong articulation point; however, the converse does not hold (e.g., node $4$ in Example 2).
-The previous discussion describes the properties of the LTP agents.
-In larger networks, however, multiple LTP agents may exist.
-This leads to the question: Can a non-influential agent be persuaded by more than one LTP agent?
+Figure 3: Node 33 is an LTP agent and it persuades 44 and $5$ as each path from the stubborn agents traverse $3$.
+Suppose a network $\mathcal{G}$ has multiple LTP agents.
+The following question naturally arises: Can an agent be persuaded by two different LTP agents?
+The following Lemma presents an answer to this question.
 Lemma 3
 Consider a network $\mathcal{G}$ and let $\mathcal{L}$ denote the set of LTP agents in $\mathcal{G}$. Then, $\mathcal{N}_{i}\cap\mathcal{N}_{j}=\emptyset$ for any distinct $i,j\in\mathcal{L}$.
-Proof 4.1.
+Proof:
 Suppose the contrary is true and there exists a node $q\in\mathcal{N}_{i}\cap\mathcal{N}_{j}$.
@@
 Similarly, we can show that a path from $s\to q$ exists that does not traverse $j$.
-Since all of the conditions: C1, C2 and C3, lead to a contradiction, the assumption $q\in\mathcal{N}_{i}\cap\mathcal{N}_{j}$ does not hold.
+Since all of the conditions: C1, C2 and C3, lead to a contradiction, it means that the assumption $q\in\mathcal{N}_{i}\cap\mathcal{N}_{j}$ does not hold.
+∎
 Lemma 3 establishes that an LTP agent and the agents it persuades form a disjoint group in the network.
-Certain agents in $\mathcal{G}$ can exist who are neither an LTP agent nor are persuaded by any LTP agent.
+However, there can be certain agents who are neither an LTP agent nor are persuaded by any LTP agent.
 Excluding them, we can uniquely partition the remaining agents into disjoint groups, with each group consisting of one LTP agent and its persuaded agents.
-4.1 The role of network topology on opinion formation
-In this subsection, we present a framework that relates the final opinions of agents with the underlying network topology.
-Subsequently, we employ this framework to demonstrate that the topological relation between an LTP agent and the agent(s) it persuades leads to the formation of an opinion cluster.
-Consider a network $\mathcal{G}$ of $n$ agents with opinions evolving under the FJ model (1).
-Let $\mathcal{G}$ have $m$ stubborn agents.
-Then, the final opinions satisfy: $\mathbf{x}^{*}=(I-\beta)W\mathbf{x}^{*}+\beta\mathbf{x}(0)$.
-We can rewrite these linear equations as follows:
-$\underbrace{\begin{bmatrix}(I-(I-\beta)W)&-\eta&\\ 0&0&\end{bmatrix}}_{R}\ \underbrace{\begin{bmatrix}\mathbf{x}^{*}\\ \mathbf{x}_{s}(0)\end{bmatrix}}_{z}=\mathbb{0}$ (2)
-where $\mathbf{x}_{s}=[x_{s_{1}}(0),...,x_{s_{m}}(0)]$ and $s_{1},s_{2},...,s_{m}\in[m]$ are the labels of the $m$ stubborn agents.
-The matrix $\eta=[\eta_{ij}]\in\mathbb{R}^{n\times m}$ has entry $\eta_{{i}j}=\beta_{i}$ if the stubborn agent $i$ is labelled as $s_{j}$, otherwise $\eta_{ij}=0$ for $i\in[n],j\in[m]$.
-Henceforth, we represent eqn.
-(2) as $Rz=\mathbb{0}$ where $R=[r_{i,j}]\in\mathbb{R}^{(n+m)\times(n+m)}$.
-Remark 4.2.
-The following are the salient properties of $R$: (a) each row sum in $R$ is 0, (b) the diagonal entries are non-negative while the off-diagonal entries are non-positive.
-Thus, $R$ is a Laplacian matrix under the Defn.
-6.3 in [20].
-The digraph $\mathcal{G}(R)$ can be induced from $R$ if an edge $(i,j)$ exists in $\mathcal{G}(R)$ when $r_{ji}<0$.
-The associated graph $\mathcal{G}(R)$ has $n+m$ nodes.
-Being consistent with the indexing of nodes in $\mathcal{G}$, the first $n$ nodes in $\mathcal{G}(R)$ are associated with $x_{i}^{*}$ for $i\in[n]$ and the nodes $n+1$ to $n+m$ are associated with the initial states of the $m$ stubborn agents.
-Each node $n+i$ forms a source in $\mathcal{G}(R)$ for $i\in[m]$.
-The source $n+i$ has only one outgoing edge to the node associated with the final opinion of the corresponding stubborn agent.
-For example, in Fig.
-3, the initial opinion of stubborn agent $2$ is associated with node $n+1$ in $\mathcal{G}(R)$, thus, $n+1$ only has outgoing edge $(n+1,2)$.
-Example 3
-Consider the network $\mathcal{G}$ in Fig.
-1(a) with stubborn agents $2$ and $6$.
-The corresponding network $\mathcal{G}(R)$, shown in Fig.
-3, is derived from eqn.
-(2).
-Each node associated with the initial opinion of a stubborn agent is shown in turquoise.
-Figure 3: The network $\mathcal{G}(R)$ derived from $\mathcal{G}$.
-Remark 4.3.
-Under Kron Reduction, we determine the Schur complement (4) of the loopy Laplacian matrix $Q$.
-Since $R$ is a Laplacian matrix, it can be reduced by Kron Reduction for a suitable choice of the subset of nodes $\alpha$.
-If $\mathcal{G}$ does not have any oblivious agents, then each node in $[n]$ has a path in $\mathcal{G}(R)$ from at least one source $\{n+1,...,n+m\}$.
-Thus, by Statement 1) in Lemma 5.1, the Schur complement $R/\alpha^{c}$ is well-defined for any $\alpha$ such that $\{n+1,...,n+m\}\subseteq\alpha\subseteq[n+m]$.
-On the contrary, if $\mathcal{G}$ has an oblivious agent, then the sources $\{n+1,...,n+m\}$ do not have a path in $G(R)$ to this agent.
-For such networks, the set $\alpha$ must include the oblivious agents belonging to each iSCC and sources $\{n+1,...,n+m\}$.
-For this choice of $\alpha$, each $i\in\alpha^{c}$ has a path in $\mathcal{G}(R)$ from a node $j\in\alpha$.
-Thus, Statement 1) in Lemma 5.1 holds.
-Consider a set of linear equations $M\mathbf{y}=\mathbb{0}$, where $\mathbf{y}\in\mathbb{R}^{p}$.
-Eliminating $y[\alpha^{c}]$ using Schur complement (4), yields the reduced system of equations: $M/\alpha^{c}\ \mathbf{y}[\alpha]=0$ (Pg.
-18 in [21]).
-Similarly, we eliminate the states $z[\alpha^{c}]$ in the linear eqns.
-(2) using the Schur Complement to obtain,
-$R/\alpha^{c}.\begin{bmatrix}\mathbf{x}^{*}[\omega]&\mathbf{x}_{s}(0)]\end{bmatrix}^{T}=\mathbb{0}$ (3)
-where $\omega=\alpha\setminus\{n+1,...,n+m\}$.
-Thus, the Kron Reduction technique is useful in determining the relations between the final opinions of a subgroup of nodes in $\mathcal{G}$ and quantifying their dependence on the influential agents.
-4.2 Main Results
-Using the Kron Reduction technique, we present topology-based conditions under which the agents in a subgroup have equal final opinions and form an opinion cluster.
-Theorem 1.
-Consider a network $\mathcal{G}$ with $m\geq 1$ stubborn agents.
-Let Assumption 1 hold.
+The following results presents the relation between the final opinions of agents within each such group.
+Theorem 1
+Consider a network $\mathcal{G}$ of $n$ agents with $m$ stubborn agents.
 The opinions of agents in $\mathcal{G}$ are governed by the FJ model (2).
-If $p$ is an LTP agent, then $p$ and the nodes in set $\mathcal{N}_{p}$ form an opinion cluster.
-The proof is given in the Appendix.
-In [15], the authors show that the final opinion of an agent depends on the paths (the weights of edges on the path) to it from the stubborn agents.
-However, it remains unclear what topological conditions on the paths from stubborn agents to distinct agents ensure that the final opinions of these agents converge to the same value.
-Theorem 1 addresses this gap by proving that the topological relation between an LTP agent and the agents it persuades leads to the formation of an opinion cluster.
-A major advantage of this result is that regardless of edge weights in the network $\mathcal{G}$ and the stubbornness values, an LTP agent $p$ and the agents in $\mathcal{N}_{p}$ collectively form an opinion cluster.
-Example 1 illustrates the result in Theorem 1 as the LTP agent $3$ and agents in $\mathcal{N}_{3}=\{4,5\}$ form an opinion cluster.
-Similarly, the LTP agent $6$ forms another opinion cluster with agent in $\mathcal{N}_{6}=\{1\}$.
-Remark 4.4.
-In [12], the authors derived the conditions to obtain $k\in\mathbb{N}$ opinion clusters in weakly connected digraphs, under the assumption that each cycle in the digraph has at most one stubborn agent.
-Theorem 1 generalises these conditions to form opinion clusters in any arbitrarily connected digraphs, thereby removing this restrictive assumption.
-Moreover, in [12] an opinion cluster is formed by group of agents composed either (1) of a single stubborn agent with the rest being non-stubborn, or (2) of all the oblivious agents in the network $\mathcal{G}$ when $\mathcal{G}$ has a unique aperiodic iSCC of oblivious agents.
-Hence, a group containing only non-influential agents does not form any opinion cluster in [12].
-In contrast, Theorem 1 establishes that an opinion cluster can be formed by any group of non-influential agents if one of them is an LTP agent $p$ and the rest are in $\mathcal{N}_{p}$ (e.g., the agents $3,4,5$ in Example 1).
-Remark 4.5.
-Consider an agent $p\in\mathcal{V}$.
-To verify if $p$ is an LTP agent, we first remove all the outgoing edges of $p$, which takes $O(n)$ time.
-Next, the modified $\mathcal{G}$ is traversed $m+t$ times using the Depth First Search (DFS), once from each of the $m$ stubborn agents and the $t$ iSCCs composed of oblivious agents.
-As one iteration of DFS has complexity $O(n+|\mathcal{E}|)$, this step requires $O((m+t)(n+|\mathcal{E}|))$ time, where $|\mathcal{E}|$ denotes the number of edges.
-If a node remains unvisited after all these traversals, then $p$ satisfies Condition (i) in Defn 2.
-Let $\mathcal{U}_{p}$ be the set of unvisited nodes.
-By repeating this step for all $p\in\mathcal{V}$, we identify all agents that satisfy Condition (i) and store them as $\mathcal{H}$.
-This step has a time complexity of $O((t+m)n(n+|\mathcal{E}|))$ which reduces to $O(n(n+|\mathcal{E}|))$ since $t+m\ll n$.
-If $\mathcal{H}\subseteq\mathcal{I}$, identifying all LTP agents requires $O(n(n+|\mathcal{E}|))$ time.
-Next, we verify condition (ii) in Defn.
-2 for each non-influential agent in $\mathcal{H}$.
-For an non-influential agent $p$, we check if there is a $c\in\mathcal{H}\setminus\{p\}$ such that $p\in\mathcal{U}_{c}$.
-This step requires $O(n^{2})$ time.
-Repeating it for all $p\in\mathcal{H}$ yields a worst-case complexity of $O(n^{3})$.
-Thus, the overall time complexity of identifying all LTP agents in $\mathcal{G}$ is $O(n^{3})$.
-Using the topological conditions in Theorems 1, a network can be designed to obtain a predefined number of opinion clusters with each cluster formed by a desired set of agents.
-Corollary 2.
-Consider a network $\mathcal{G}$ with $m\geq 1$ stubborn agents and $t\geq 0$ iSCCs with oblivious agents such that Assumption 1 holds.
-The opinions of agents in $\mathcal{G}$ are governed by the FJ model (1).
-Let $\mathcal{O}$ be the set of the oblivious agents that either belong to an iSCC or have paths from only one iSCC in $\mathcal{G}$.
-Then, $k\geq t+m$ opinion clusters form in $\mathbf{x}^{*}$ if the agents in set $\mathcal{V}\setminus\mathcal{O}$ can be partitioned into ktk-t disjoint groups with each group composed of an LTP agent and the nodes persuaded by it.
-The proof of Corollary 2 is given in the Appendix. The following example illustrates the conditions presented in Corollary 2 to obtain the desired opinion clustering.
+If each iSCC in $\mathcal{G}$ composed of the oblivious agents is aperiodic, then final opinions of an LTP agent $p$ and nodes in set $\mathcal{N}_{p}$ form an opinion cluster.
+The proof of Theorem 1 is given in the Appendix.
+Theorem 1 establishes that an LTP agent $p$ and the agents in $\mathcal{N}_{p}$ collectively form an opinion cluster in the final opinion.
+Note that the existence of an LTP agent in a network depends only on paths from the influential agents; hence, it is a purely graph topological property.
+Consequently, an LTP agent $p$ and $\mathcal{N}_{p}$ form an opinion cluster independent of the edge weights and the stubbornness of the stubborn agents.
+Example 1 illustrates the result in Theorem 1 as the LTP agent $3$ forms an opinion cluster with agents in $\mathcal{N}_{3}=\{4,5\}$ and LTP agent $6$ forms another opinion cluster with agent in $\mathcal{N}_{6}=\{1\}$.
+Additionally, it explains the formation of a new opinion cluster for the modified graph $\mathcal{\hat{G}}$.
+This occurs because the addition of edge $(1,5)$ results in a path from $6$ to $5$ that does not traverse $3$.
+Thus, $5\notin\mathcal{N}_{3}$ and it does not form an opinion cluster with $3$ and $4$, as shown in Fig.
+1(d).
+By definition, if an LTP agent $p$ is oblivious, then the agents in $\mathcal{N}_{p}$ are also oblivious.
+Consequently, it follows from Theorem 1 that a set of non-influential oblivious agents can also form an opinion cluster.
+Since the opinion of an oblivious agent evolves according to DeGroot model and is not affected by stubborn behaviour, an LTP agent ensures the formation of opinion clusters even in DeGroot framework.
+Note that since nodes in $\mathcal{N}_{q}$ are always non-influential, the desired opinion clusters can be formed in DeGroot framework only in weakly connected networks.
+Remark 3
+In [15], a network with at most one stubborn agent in each cycle is shown to form opinion clusters under the FJ model.
+Theorem 1 generalises the existing results by presenting LTP-based conditions to form an opinion cluster in any arbitrary digraph.
+Further, in [15], each opinion cluster is formed by a group that contains at least one influential agent: a stubborn agent with non-stubborn agents or the oblivious agents in an iSCC.
+In contrast, Theorem 1 demonstrates that a group composed only of non-influential agents can also form an opinion cluster provided that one of them is an LTP agent and the rest are persuaded by it.
+Next, we illustrate the topological conditions presented in Theorem 1 using the following example.
 Example 4
@@
 4(b).
-One cluster is formed by the oblivious agent belonging to the iSCC and the rest are formed by the agents belonging to the same group.
+One cluster is formed by the oblivious agents in the iSCC and the rest are formed by the $4$ disjoint groups of agents; each composed of an LTP agent and those persuaded by it.
 (a) Network $\mathcal{G}$
@@
 Figure 4: Suitable design of $\mathcal{G}$ for desired opinion clusters.
-5 Conclusions
+VI Conclusions
 The FJ model incorporates the individual biases as stubborn behaviour in the averaging-based opinion models.
@@
 Every path from an influential agent (a stubborn agent or an oblivious agent in an iSCC) to the persuaded agent always traverses the LTP agent.
-Using Kron Reduction, we establish in Theorem 1 that an LTP agent and the set of agents it persuades form an opinion cluster in any arbitrarily-connected digraph under the FJ framework.
-Some additional key insights from our results are as follows: (i) The presence of LTP agents always results in opinion clusters, independent of the edge-weights in the network and the stubbornness of the agents.
-(ii) The notion of LTP agents generalises the topology-based conditions for opinion clustering in [12] to any arbitrary digraphs.
-(iii) By suitably placing the LTP agents and designing the network topology, we achieve any desired opinion clustering.
-In future, we plan to use the LTP agents for desirably shaping the final opinions in real-world social networks.
-Appendix
-Kron Reduction: Let $M\in\mathbb{R}^{p\times p}$ with index set $\alpha\subseteq[p]$ and $\alpha^{c}:=[p]\setminus\alpha$.
-If $M[\alpha^{c}]$ is non-singular, then the Schur complement of $M[\alpha^{c}]$ in $M$ is defined as:
-$M/\alpha^{c}=M[\alpha]-M[\alpha,\alpha^{c}](M[\alpha^{c}])^{-1}M[\alpha^{c},\alpha]$ (4)
-Under Kron Reduction, the Schur complement of the loopy Laplacian matrix $Q$ is evaluated.
-Lemma 5.1 ([18]).
-Consider a loopy Laplacian matrix $Q\in\mathbb{R}^{p}$ with $\alpha\subset[p]$. Let $\mathcal{G}(Q)$ be the digraph associated with $Q$.
-  1. 1.
-    $Q/\alpha^{c}$ is well defined ( i.e.
-$Q[\alpha^{c}]$ is invertible) if for each node $i\in\alpha^{c}$ there is a node $j\in\alpha$ such that a path $j\to i$ exists in $\mathcal{G}(Q)$.
-111(Note that we adapt this condition to our framework because we consider $w_{ij}>0$ when $(j,i)\in\mathcal{E}$.)
-  2. 2.
-    If $Q$ is a loopless Laplacian matrix, then $Q/\alpha^{c}$ is also a loopless Laplacian matrix.
-5.1 Proof of Theorem-1
-Before moving to the proof of Theorem-1, the following Lemma needs to be stated.
-Lemma 5.2.
+Using Kron reduction, we establish in Theorem 1 that the final opinions of an LTP agent and the set of agents it persuades are equal.
+Thus, they form an opinion cluster in any arbitrarily-connected digraph under the FJ framework.
+Interestingly, such LTP agents lead to the formation of opinion clusters even under the DeGroot model (see Theorem 1).
+Some additional key insights from our results are: (i) The presence of LTP agents always results in opinion clusters, independent of the edge-weights in the network and the stubbornness of the agents.
+(ii) Notably, an LTP agent may or may not be influential, but still shapes the final opinions of the agents it persuades.
+(iii) The notion of LTP agents generalises the topology-based conditions for opinion clustering in [15] to any arbitrary digraphs.
+(iv) By suitably placing the LTP agents and designing the network topology, we achieve any desired opinion clustering as demonstrated in Example 4.
+In future, we plan to examine the potential of LTP agents in desirably shaping the final opinions in social networks.
+VII Appendix
+Before moving to the proof of Theorem 1, the following Lemma needs to be stated.
+Lemma 4
 Consider a weakly connected network $\mathcal{G}$ with $m\geq 1$ stubborn agents and matrix $R$ derived using eqn.
 (3).
-Let $p$ be an LTP agent and $q\in\mathcal{N}_{p}$.
-If $\alpha$ is the union of set $\{p,q,n+1,....,n+m\}$ and all the oblivious agents in iSCC in $\mathcal{G}$ and $\alpha^{c}=[n+m]\setminus\alpha$ such that $\alpha^{c}\neq\phi$.
+Let $\alpha$ be a set containing $\{n+1,....,n+m\}$ and all the oblivious agents in the iSCCs in $\mathcal{G}$ and $\alpha^{c}=[n+m]\setminus\alpha$ such that $\alpha^{c}\neq\phi$.
 Then,
 $(R[\alpha^{c}])^{-1}=\sum_{k=0}^{\infty}\big((I-\beta[\alpha^{c}])W[\alpha^{c}]\big)^{k}$ (5)
-Proof 5.3.
-By definition of $\alpha$, we get $\alpha^{c}\subset[n]$.
+Proof:
+By definition of $\alpha$, we get $\alpha^{c}\subseteq[n]$.
 From the construction of $R$, it follows that $R[\alpha^{c}]=I-(I-\beta[\alpha^{c}])W[\alpha^{c}]$.
-If $\rho((I-\beta[\alpha^{c}])W[\alpha^{c}])<1$, then identity (5) follows by Neumann series.
 Let $M=(I-\beta[\alpha^{c}])W[\alpha^{c}]$.
-To prove that $\rho(M)<1$, it suffices to show that (I) $M$ is row substochastic, (II) for each $i^{th}$ row in $M$ with row-sum is $1$ there exists a $j^{th}$ row with row-sum less than $1$ such that there is a path from $j$ to $i$ in the associated digraph $\mathcal{G}(M)$.
-Then, by Theorem 4.11 in [20])1, $\rho(M)<1$.
-Step-I: $M$ is row-substochastic if: (i) a node $i\in\alpha^{c}$ whose in-neighbour is $p,q$ or an oblivious agent in iSCC satisfies $\sum_{j\in\alpha^{c}}w_{ij}<1$, (ii) node $i\in\alpha^{c}$ is a stubborn agent and the corresponding row-sum is at most $(1-\beta_{i})<1$.
-In an arbitrary digraph $\mathcal{G}$, the following scenarios may occur:
-  • $h\in\alpha^{c}$ has an oblivious agent: Due to weak connectivity of $\mathcal{G}$, an oblivious agent in iSCC exists in $\alpha$ that has a path to $h$, implying that (i) holds.
-  • All agents in $\alpha^{c}$ are non-oblivious and a stubborn agent exists in $\alpha^{c}$: in this case, (ii) holds for the stubborn agent.
-  • All agents in $\alpha^{c}$ are non-oblivious and non-stubborn: Since $q$ cannot be stubborn, node $p$ is the only stubborn agent in $\mathcal{G}$.
-Thus, at least one node exists in $\alpha^{c}$ that has either $p$ or $q$ as in-neighbour, implying that (i) holds.
-Thus, there always exists a node in $\alpha^{c}$ that satisfies either (i) or (ii). Hence, $M$ is row substochastic.
-Step-II: As discussed in Step-I, each oblivious agent has a path from an oblivious agent in iSCC in $\mathcal{G}$.
-It follows from (i) that this path traverses an oblivious agent in $\alpha^{c}$ whose row-sum in $M$ is less than $1$.
-A special case could occur if each path in $\mathcal{G}$ from the oblivious agents in iSCC to an oblivious agent $\tilde{j}\in\alpha^{c}$ either traverses $p$ or $q$.
-In this case, the in-neighbour(s) of $p$ or $q$ in $\alpha^{c}$ who has row-sum in $M$ less than $1$ lie on this path.
-Hence, every oblivious agent in $\alpha^{c}$ with row sum $1$ has a path in $\mathcal{G}(M)$ from a node with row sum <1<1.
-A similar argument applies to the non-oblivious agents as well.
-Thus, statement II also holds and $\rho(M)<1$.
-Proof of Theorem-1: Let $q\in\mathcal{N}_{p}$.
+Since $W$ is row-stochastic, $M$ satisfies $M\mathbb{1}_{n}=(I-\beta)\mathbb{1}_{n}\leq\mathbb{1}_{n}$.
+Hence, the row-sum of any row in $M$ can be at most $1$.
+By the Gersgorin disk theorem, the spectrum of $M$ is a subset of a unit disk in the complex plane (Theorem 2.9 in [17]).
+Further, since $M$ is a non-negative matrix, by the Perron-Frobenius Theorem, there exists a real eigenvalue $\lambda_{M}$ such that $\lambda_{M}\geq|\mu|\geq 0$, where $\mu$ are the remaining eigenvalues of $M$ (Theorem 2.12 in [17]).
+Thus, $\lambda_{M}\in[0,1]$.
+Further, as discussed in Remark 2, the Schur complement is well defined for any $\alpha$ containing $\{n+1,....,n+m\}$ and the oblivious agents in iSCCs of $\mathcal{G}$.
+Thus, $R[\alpha^{c}]$ is invertible.
+This means that none of the eigenvalues of $M$ is equal to $1$.
+Thus, $\lambda_{M}<1$ and eqn.
+(5) follows by Neumann series.
+∎
+Proof of Theorem 1: Let $q\in\mathcal{N}_{p}$.
 To prove that the final opinion $x_{p}^{*}=x_{q}^{*}$, we reduce the matrix $R$ in eqn.
@@
 Consider $\alpha$ to be $\{p,q,1,...,u,n+1,...,n+m\}$.
-From Remark 4.3, we know that the Kron Reduction $R/\alpha^{c}$ is well-defined.
+From Remark 2, we know that the Kron Reduction $R/\alpha^{c}$ is well-defined.
 Next, we evaluate $R/\alpha^{c}$ given by eqn.
@@
 Thus, the corresponding rows of $R$ have all entries as zero.
-Therefore, $R[\alpha]$ takes the following form:
-$R[\alpha]={\begin{bmatrix}r_{p,p}&r_{p,q}&R[p,[u]]&r_{p,n+1}&...&r_{p,n+m}\\ r_{q,p}&r_{q,q}&\mathbb{0}&0&...&0&\\ \mathbb{0}&\mathbb{0}&R[[u]]&\mathbb{0}&\cdots&\mathbb{0}\\ \mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&...&\mathbb{0}_{m\times 1}\end{bmatrix}}$.
-Next, we consider $Y=R[\alpha,\alpha^{c}](R[\alpha^{c}])^{-1}R[\alpha^{c},\alpha]$. From Lemma 5.2, it follows that
+Therefore, $R[\alpha]$ takes the following form: $\begin{bmatrix}r_{p,p}&r_{p,q}&R[p,[u]]&r_{p,n+1}&...&r_{p,n+m}\\ r_{q,p}&r_{q,q}&\mathbb{0}&0&...&0&\\ \mathbb{0}&\mathbb{0}&R[[u]]&\mathbb{0}&\cdots&\mathbb{0}\\ \mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&...&\mathbb{0}_{m\times 1}\end{bmatrix}$}.
+Next, we consider $Y=R[\alpha,\alpha^{c}](R[\alpha^{c}])^{-1}R[\alpha^{c},\alpha]$. From Lemma 4, it follows that
 $Y=R[\alpha,\alpha^{c}]\sum_{k=0}^{\infty}\big((I-\beta[\alpha^{c}])W[\alpha^{c}]\big)^{k}R[\alpha^{c},\alpha]$ (6)
-Let $F:[{u}+m+2]\to\alpha$ be a mapping such that $F(1)=p,F(2)=q$, $F(j+2)=j$ for $j\in[u]$, and $F(u+2+j)=n+j$ for $j\in[m]$.
-This mapping relates the row (and column) indices of $Y$ (and $R/\alpha^{c}$) with the agents in set $\alpha$.
+Let $F:\alpha\to[{u}+m+2]$ be a mapping such that $F(p)=1,F(q)=2$, $F(j)=j+2$ for $j\in[u]$, and $F(n+j)=u+2+j$ for $j\in[m]$.
+This mapping relates the agents in set $\alpha$ with the row (and column) indices of $Y$ (and $R/\alpha^{c}$).
 Then, from eqn.
-(6), the entry $y_{ij}$ of $Y$ is non-zero only if there exists a walk in $\mathcal{G}(R)$ from $F(j)$ to $F(i)$ such that each node on this walk (except $F(i)$ and $F(j)$) belongs to $\alpha^{c}$.
-By property P1), $y_{2j}=0$ for all $j\in{\{3,...,u+m+2\}}$ because each walk from corresponding $F(j)$ to $q$ in $\mathcal{G}(R)$ always passes through $p\in\alpha$.
+(6), the entry $y_{F(i)F(j)}$ of $Y$ is non-zero only if there exists a walk in $\mathcal{G}(R)$ from $j$ to $i$ such that each node on this walk (except $i$ and $j$) belongs to $\alpha^{c}$.
+By property P1), $y_{2F(j)}=0$ for all $j\in[u]\cup\{n+1,...,n+m\}$ because each walk from corresponding $j$ to $q$ in $\mathcal{G}(R)$ always passes through $p\in\alpha$.
 Consequently, $R/\alpha^{c}=R[\alpha]-Y$ has the form:
-$\begin{bmatrix}r_{1,1}^{1}&r_{1,2}^{1}&R_{1}&r_{1,u+2}^{1}&...&r_{1,u+m+2}^{1}\\ r_{2,1}^{1}&r_{2,2}^{1}&\mathbb{0}_{u}&0&...&0\\ \mathbb{0}_{u\times 1}&\mathbb{0}_{u\times 1}&R[[u]]&\mathbb{0}_{u\times 1}&\cdots&\mathbb{0}_{u\times 1}\\ \mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&...&\mathbb{0}_{m\times 1}\end{bmatrix}$}
-where $r_{i,j}^{1}$ gives the entry of matrix $R/\alpha^{c}$ for $i,j\in[m+2]$.
-The matrix $R_{1}$ is the submatrix $R/\alpha^{c}[1,[u]]$ highlighting the corresponding non-zero entries.
-From Lemma 5.1, we know that $R/\alpha^{c}$ is also a Laplacian matrix and the row-sum of each of its rows equals 0.
+$\begin{bmatrix}r_{1,1}^{1}&r_{1,2}^{1}&\Delta_{1}&r_{1,u+3}^{1}&...&r_{1,u+m+2}^{1}\\ r_{2,1}^{1}&r_{2,2}^{1}&\mathbb{0}_{u\times 1}&0&...&0\\ \mathbb{0}_{u\times 1}&\mathbb{0}_{u\times 1}&R[[u]]&\mathbb{0}_{u\times 1}&\cdots&\mathbb{0}_{u\times 1}\\ \mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&...&\mathbb{0}_{m\times 1}\end{bmatrix}$
+where $r_{i,j}^{1}$ gives the entry of matrix $R/\alpha^{c}$ for $i,j\in[m+u+2]$.
+The vector $\Delta_{1}=[r_{1,3}^{1},r_{1,4}^{1},...,r_{1,u+2}^{1}]$ highlights the corresponding non-zero entries.
+From Lemma 1, we know that $R/\alpha^{c}$ is also a Laplacian matrix and the row-sum of each of its rows equals 0.
 Since the second row of $R/\alpha^{c}$ has only two non-zero entries, thus, $r_{2,1}^{1}=-r_{2,2}^{1}$.
-Hence, by eliminating the states $\alpha^{c}$ in the steady state eqns.
+Moreover, by eliminating the states $\alpha^{c}$ in the steady state eqns.
 (3), we get $x_{p}^{*}=x_{q}^{*}$.
 Since $q$ is any node in $\mathcal{N}_{p}$, this holds for all nodes in $\mathcal{N}_{p}$.
-$\blacksquare$
-5.2 Proof of Corollary 2
-We classified the agents in $\mathcal{G}$ into sets $\mathcal{O}$ and $\mathcal{V}\setminus\mathcal{O}$.
-As $\mathcal{V}\setminus\mathcal{O}$ is union of tkt-k disjoint sets $\{i\}\cup\mathcal{N}_{i}$, where each $i\in\mathcal{L}$ is an LTP agent.
-Hence, ktk-t LTP agents are present in set $\mathcal{V}\setminus\mathcal{O}$.
-By Theorem 1, each LTP agent $i$ and the agents in $\mathcal{N}_{i}$ form an opinion cluster.
-Therefore, the agents in set $\mathcal{V}\setminus\mathcal{O}$ form ktk-t clusters.
-The set $\mathcal{V}\setminus\mathcal{O}$ contains the $m$ stubborn agents.
-Since stubborn agents persuaded by an LTP agent, each stubborn agent must be an LTP agent.
-Thus, $k-t\geq m$.
-[…]
-Let $C_{j}$ be the set of oblivious agents in $j^{th}$ iSCC and those oblivious agents only have a path from the $j^{th}$ iSCC.
-The opinions of agents in $C_{j}$ evolve as $\mathbf{y}_{j}(k)=W_{i}^{j}\mathbf{y}_{j}(k)$ with $\mathbf{y}_{j}$ is the vector that denotes the opinion of agents in this set.
-Since the associated digraph $G(W_{i}^{t})$ is a subgraph of $\mathcal{G}$ containing a unique aperiodic iSCC, the opinions of the agents converge to consensus (Theorem 5.1 in [20])1.
-Hence, the $t$ subgroups, each composed of an iSCC of oblivious agents and those having a path only from the iSCC, form $t$ opinion clusters.
-Collectively, the oblivious agents in $\mathcal{O}$ and the agents in $\mathcal{V}\setminus\mathcal{O}$ form $k$ opinion clusters such that $k\geq t+m$.
+$\blacksquare$.